### v1

# An ordering for the strength of functional dependence

Abstract
We introduce a new dependence order that satisfies eight natural axioms that we propose for a global dependence order. Its minimal and maximal elements characterize independence and perfect dependence. Moreover, it characterizes conditional independence, satisfies information monotonicity, and exhibits several invariance properties. Consequently, it is an ordering for the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}\,.$ As we show, various dependence measures, such as Chatterjee’s rank correlation, are increasing in this order. We characterize our ordering by the Schur order and by the concordance order, and we verify it in models such as the additive error model, the multivariate normal distribution, and various copula-based models.
Keywords Chatterjee’s rank correlatifon, concordance order conditional independence, convex order, copula, dependence measure, dimension reduction, information gain inequality, rearrangements, perfect dependence, Schur order
1 Introduction and main results
In recent years, a large body of statistical literature has focused on dependence measures taking values in $[0,1]$, where 0 characterizes independence, and $1$ perfect functional dependence. A prominent example is Chatterjee’s rank correlation [7, 15] defined in Equation (2) below. While this measure gained popularity due to its fast nearest-neighbor–based estimator, it is also theoretically appealing due to further remarkable properties: it satisfies information monotonicity and characterizes conditional independence, enabling important applications such as model-free variable selection; see [16] for a survey. An open question concerns the identification of a dependence order that underlies the above type of dependence measures and describes their behavior across various models. In this paper, we fill this gap by introducing a new dependence order that allows us to answer the following question in the affirmative.
Is there a dependence order that characterizes independence through its minimal elements and perfect functional dependence through its maximal elements, satisfies information monotonicity and characterizes conditional independence, underlies Chatterjee’s rank correlation and related dependence measures, has a simple and interpretable form, can be verified in a variety of models, and applies to a large class of distributions?
Classical rank correlations such as Kendall’s $\tau$, Spearman’s $\varrho$, or Gini’s $\gamma$ quantify the degree of positive or negative dependence among random variables. There is a large literature on positive dependence orderings underlying these rank correlations; see [9, 24, 31, 35, 36, 37, 55, 56] and the references therein. The certainly most common positive dependence order is the concordance order [63, 68], defined by a pointwise comparison of distribution and survival functions; see (27). Its minimal and maximal elements correspond to countermonotonocity and comonotonocity, i.e., to perfect negative and positive monotone dependence, respectively.
In contrast, for measures of regression dependence—which have become highly popular over the last decade—there is relatively little literature on underlying dependence orderings. Apart from our preliminary work [3, 5], research on global dependence orderings is conducted in [20, 22, 30, 54, 58, 59]. However, none of the dependence orderings considered in these references is able to describe, even in the simplest models, the behavior of dependence measures that characterize both independence and perfect functional dependence. For example, consider the additive error model
$Y=f(\mathbb{X})+\sigma\varepsilon,$ (1)
where $\sigma\geq 0$ is a deterministic parameter scaling the noise $\varepsilon$ that is assumed to be independent of $f(\mathbb{X})\,.$ Of course, a suitable dependence measure quantifying the strength of dependence of $Y$ on $\mathbb{X}$ should be decreasing in $\sigma.$ Intuitively, one would attribute such a property to Chatterjee’s $\xi$; however, to the best of our knowledge, such ordering results for $\xi$ are known only under specific model assumptions, such as convolution semigroups. In the latter case, the desired monotonicity of $\xi$ in $\sigma$ is a direct consequence of the data processing inequality in [17]; see Remark 1.9(b) below.
In this paper, we introduce a new global dependence order—the conditional convex order—that describes the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}$. This ordering reflects the fundamental properties of a large class of dependence measures including $\xi$, and we verify it in settings like the additive error model (1), the multivariate normal distribution, and various copula-based models. Before we propose eight natural axioms for a global dependence ordering underlying $\xi$, recall the population version of Chatterjee’s rank correlation defined by
$\xi(Y,\mathbb{X}):=\frac{\int_{\mathbb{R}}\mathrm{Var}(P(Y\geq y\,|\,\mathbb{X}))\;\mathrm{d}P^{Y}(y)}{\int_{\mathbb{R}}\mathrm{Var}(\mathds{1}_{\{Y\geq y\}})\;\mathrm{d}P^{Y}(y)};$ (2)
see [7, 15]. It relies on the copula-based version in [22] and on the sensitivity measures in [12, 26]. Dependence measures like $\xi$ capture the variability of conditional distributions in various ways and have attracted a lot of attention in the past few years; see e.g. [6, 10, 13, 14, 21, 28, 29, 43, 60, 62, 67]. They attain values in the interval $[0,1]$ and have the zero-independence and max-functionality property, i.e., the minimal value 0 characterizes independence of $\mathbb{X}$ and $Y$, while the maximal value $1$ characterizes perfect (functional) dependence of $Y$ on $\mathbb{X}\,.$ The latter means that there exists a measurable function $f$ (not necessarily increasing or decreasing) such that $Y=f(\mathbb{X})$ almost surely.
For understanding the behavior of dependence measures in models such as (1), a suitable dependence order is essential. It is straightforward to see that $\xi$ is not increasing in the concordance order. For example, let $X$ and $Y$ be standard normal with $Y=\pm X.$ Then $(Y,X)$ is comonotone/countermonotone and thus maximal/minimal in the concordance order, but $\xi(Y,X)=1$ in both cases since $Y$ is a deterministic function of X.X\,.
1.1 Axioms for orderings of global dependence
We propose the following axioms for a global dependence ordering $\prec$ that compares the strength of functional dependence of $Y$ on $\mathbb{X}$ on a class $\mathcal{R}$ of random vectors. Well known axioms for positive dependence orderings are studied in [36, 37]. We denote by $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ equality in distribution and write $\equiv$ if $\prec$ and $\succ$ holds true.
  1. (O1)
    Law-invariance: $(Y,\mathbb{X})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y^{\prime},\mathbb{X}^{\prime})$ implies $(Y,\mathbb{X})\equiv(Y^{\prime},\mathbb{X}^{\prime})$.
  2. (O2)
    Quasi-ordering: $\prec$ is reflexive and transitive.
  3. (O3)
    Characterization of independence: $(Y,\mathbb{X})\prec(Y^{\prime},\mathbb{X}^{\prime})$ for all $(Y^{\prime},\mathbb{X}^{\prime})$ if and only if $\mathbb{X}$ and $Y$ are independent.
  4. (O4)
    Characterization of perfect dependence: $(Y^{\prime},\mathbb{X}^{\prime})\prec(Y,\mathbb{X})$ for all $(Y^{\prime},\mathbb{X}^{\prime})$ if and only if $Y$ is perfectly dependent on $\mathbb{X}\,.$
  5. (O5)
    Information monotonicity: $(Y,\mathbb{X})\prec(Y,(\mathbb{X},\mathbb{Z}))$ for all $\mathbb{X},\mathbb{Z}$ and $Y$.
  6. (O6)
    Characterization of conditional independence: $(Y,\mathbb{X})\equiv(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}\,.$
  7. (O7)
    Transformation invariance: $(Y,\mathbb{X})\equiv(g(Y),{\bf h}(\mathbb{X}))$ for every strictly increasing function $g$ and for every bijective function $\bf h$}\,.
  8. (O8)
    Distributional invariance: $(Y,\mathbb{X})\equiv(F_{Y}(Y),\mathbb{X})$ and $(Y,\mathbb{X})\equiv(Y,{\bf F}_{\mathbb{X}}(\mathbb{X}))$, where ${\bf F}_{\mathbb{X}}=(F_{X_{1}},F_{X_{2}},\ldots)$ for $\mathbb{X}=(X_{1},X_{2},\ldots)$.
Remark 1.1
  1. (a)
    While $Y$ and $Y^{\prime}$ are univariate random variables, $\mathbb{X}$ and $\mathbb{X}^{\prime}$ are random vectors that may have arbitrary dimensions $p,p^{\prime}\in\mathbb{N}$, which are not assumed to be equal.
  2. (b)
    Axiom 1 states that $\prec$ is invariant under all versions of $(Y,\mathbb{X})$. Note that the term law-invariance is also used in the context of risk measures [25].
  3. (c)
    Axiom 2 on reflexivity and transitivity is also considered in [20, 22, 58]. In contrast to positive dependence orderings, a suitable global dependence ordering $\prec$ cannot be expected to be antisymmetric (in the sense that $(Y,\mathbb{X})\equiv(Y^{\prime},\mathbb{X}^{\prime})$ implies $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$ have equal laws) because it is desirable that $(Y,Y)\equiv(Y,-Y)$ due to 4.
  4. (d)
    Axioms 3 and 4 state that independent and perfectly dependent random vectors are minimal and maximal elements over the whole class $\mathcal{R}$, i.e., they are globally extreme elements. These axioms are significantly stronger than
    • [58, Axioms O4 and O5], according to which any bivariate random vector $(Y,X)$ is comparable with an independent version $(Y^{\perp},X^{\perp})$ and with a perfectly dependent version [^1] For a bivariate random vector $(Y,X)$, an independent resp. perfectly dependent version is a random vector $(Y^{\perp},X^{\perp})$ resp. $(Y^{\top},X^{\top})$ with $Y^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\top}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $X^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\top}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X$ such that $Y^{\top}$ and $X^{\top}$ are independent resp. $Y^{\top}$ is perfectly dependent on $X^{\top}$. An independent version exists if the underlying probability space is rich enough, for example, non-atomic. A perfectly dependent version exists if $\overline{\mathsf{Ran}(F_{Y})}\subseteq\overline{\mathsf{Ran}(F_{X})}$, where $\mathsf{Ran}(F_{Z})$ denotes the range of the distribution function of $Z$, and $\overline{A}$ is the closure of a set $A\subseteq[0,1]$. $(Y^{\top},X^{\top})$ from the same Fréchet class;
    • [22, Axiom O2 and O3], according to which $(Y^{\top},X^{\top})\prec(Y,X)$ implies that $Y$ perfectly depends on $X$, and $(Y,X)\prec(Y^{\perp},X^{\perp})$ implies that $X$ and $Y$ are independent;
    • [20, Axioms O4 and O5] where minimal and maximal elements satisfy $\mathbb{E}[Y|\mathbb{X}]=\mathbb{E}Y$ a.s. and $\mathbb{E}[Y|\mathbb{X}]=Y$ a.s., respectively.
  5. (e)
    Axioms 5 and 6 state that additional information provided by further predictor variables increases the strength of functional dependence, and remains unchanged precisely when the additional information is irrelevant given the existing information. For dependence measures, information monotonicity is studied in [13]. On the level of dependence ordering, both axioms are new to the best of our knowledge. They yield a data processing inequality for $\prec$ in the sense that $(Y,\mathbb{Z})\prec(Y,\mathbb{X})$ whenever $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. This implies $(Y,{\bf h}(\mathbb{X}))\prec(Y,\mathbb{X})$ for all measurable functions $\bf h$}.
  6. (f)
    Axiom 7 is stronger than the invariance properties in [20, Axiom O3] and [59, Axiom O3] where the function $g$ is assumed to be (increasing) linear. In [22], distributions from the same Fréchet class are transformed to the copula setting and no further invariance properties are studied.
  7. (g)
    Axiom 8 on distributional invariance implies that $(Y,\mathbb{X})\prec(Y^{\prime},\mathbb{X}^{\prime})$ only depends on the underlying copulas [^2] A $d$-variate copula is the distribution function of a random vector $(U_{1},\ldots,U_{d})$ with $U_{i}$ uniform on $(0,1)$ for all $i$. By Sklar’s theorem, every $d$-variate distribution function $F$ can be decomposed into $F=C\circ(F_{1},\ldots,F_{d}),$ where $F_{i}$ is the $i$th marginal distribution function and $C$ is a copula that is unique on $\mathsf{Ran}(F_{1})\times\cdots\times\mathsf{Ran}(F_{d})$. $C_{Y,\mathbb{X}}$ and $C_{Y^{\prime},\mathbb{X}^{\prime}}$ evaluated on the Cartesian products $\mathsf{Ran}(F_{Y})\times\mathsf{Ran}(F_{X_{1}})\times\cdots\mathsf{Ran}(F_{X_{p}})$ and $\mathsf{Ran}(F_{Y^{\prime}})\times\mathsf{Ran}(F_{X_{1}^{\prime}})\times\cdots\mathsf{Ran}(F_{X_{p^{\prime}}^{\prime}})$ of the respective marginal distribution functions. More specifically, since copulas are continuous, $(Y,\mathbb{X})\prec(Y^{\prime},\mathbb{X}^{\prime})$ only depends on the closed Cartesian products.
1.2 A new global dependence order
To explain the idea how to construct a suitable dependence order that satisfies all the axioms above, let us have a closer look at the following representation of Chatterjee’s rank correlation:
$\xi(Y,\mathbb{X})=\alpha\,\int_{\mathbb{R}^{p+1}}P(Y\geq y\mid\mathbb{X}=\mathbb{x})^{2}\;\mathrm{d}(P^{\mathbb{X}}\otimes P^{Y})(\mathbb{x},y)-\beta,$ (3)
where $\alpha$ and $\beta$ are positive constants depending only on $\overline{\mathsf{Ran}(F_{Y})}$. Note that $Y$ and $\mathbb{X}$ are generally dependent while the measure integrated against is the product measure as a consequence of the definition of $\xi$. It is straightforward to see that a pointwise ordering of the integrand in (3) is too strong to define a suitable dependence order. If we split the integral in (3) into a double integral, there are two natural approaches for an ordering of conditional survival functions in squared mean.
The first approach considers the variability of the conditional survival function $P(Y\geq y\mid\mathbb{X}=\mathbb{x})$ in $y$ for each fixed $\mathbb{x}\,.$ This might be seen as the more natural and more intuitive approach supported by a variety of literature on stochastic orderings for conditional distributions; see e.g. [11, 39, 51, 65, 66, 69]. In the setting of bivariate copulas, this approach is studied in [22, Theorem 2] where the dilation order and the dispersive order are used to construct regression dependence orders, which imply an ordering of $\xi$. However, as we discuss in Section 5.2, these orders are limited because they cannot be verified even in the simplest settings.
The second approach, which we explore in this paper, examines the variability of $P(Y\geq y\mid\mathbb{X}=\mathbb{x})$ in $\mathbb{x}$ for each fixed y.y\,. This perspective relates to the theory of random measures [34], and the works closest to our paper are [58, 59] where the variability of regression functions is studied. As it turns out, our approach is much more powerful as it yields an ordering that satisfies all the axioms above and thus reflects the fundamental properties of $\xi$. Further, as we show in Theorem 1.8 and Section 3, this order can be verified for many standard models, including the additive error model (1), the multivariate normal distribution, and various copula-based models.
Pursuing the second approach, we introduce below a new dependence order that we denote as conditional convex order. It compares conditional survival probabilities with respect to the convex order—the most common variability order studied in the literature; see [57, Section 3]. To this end, recall that, for bounded random variables $S$ and $T$, the convex order $S\leq_{cx}T$ is defined by $\mathbb{E}\varphi(S)\leq\mathbb{E}\varphi(T)$ for all convex functions $\varphi\colon\mathbb{R}\to\mathbb{R}$. We denote by $q_{Y}(v):=\inf\{y\mid F_{Y}(y)\geq v\},$ $v\in(0,1)$, the left-continuous quantile function of $Y$, also known as generalized inverse of $F_{Y}$.
Definition 1.2 (Conditional convex order)
Let $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$ be random vectors. Then we define the conditional convex order $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ by $P(Y\geq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\geq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for all $v\in(0,1)$.
We write $(Y,\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ if $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ and $(Y,\mathbb{X})\succcurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$.
Before we present our main results, let us discuss why the conditional convex order is a suitable candidate for a global dependence order underlying $\xi$. First note that a necessary condition for $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. This follows from the definition of the conditional convex order using that $S\leq_{cx}T$ implies $\mathbb{E}S=\mathbb{E}T$. We refer to this condition as marginal constraint. In this case, the denominator of $\xi$ in (2) is equal for $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$, and the conditional convex order only refers to the respective numerators.
Second, Chatterjee’s rank correlation is increasing in the conditional convex order: Since $S\leq_{cx}T$ implies $\mathrm{Var}(S)\leq\mathrm{Var}(T)$, it can easily be verified that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ yields $\xi(Y,\mathbb{X})\leq\xi(Y^{\prime},\mathbb{X}^{\prime})$. We provide this statement in a more general framework in Theorem 4.1.
Third, as studied in [13], convexity is essential for information monotonicity and, in particular, for the zero-independence and max-functionality property of dependence measures. It is straightforward to verify that minimal and maximal elements in conditional convex order characterize zero-independence and max-functionality: $Y$ and $\mathbb{X}$ are independent if and only if, for all $v\in(0,1)$, $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ is a.s. constant and thus minimal in convex order. Similarly, $Y$ perfectly depends on $\mathbb{X}$ if and only if, for all $v\in(0,1)$, $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ attains a.s. values in $\{0,1\}$ and thus is maximal in convex order within the class of $[0,1]$-valued random variables with mean $P(Y\geq q_{Y}(v))$. More generally, the conditional convex order allows a simple interpretation: little/large variability of the conditional probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order indicates low/strong dependence of $Y$ on $\mathbb{X}$—with the extreme cases of independence and perfect dependence discussed before.
Remark 1.3
  1. (a)
    To derive ordering results for models such as (1), we allow in Definition 1.2 a comparison of random vectors with different marginal distributions. If $Y$ and $Y^{\prime}$ have the same distribution, $(Y,\mathbb{X}),\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ simplifies to $P(Y\geq y\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\geq y\mid\mathbb{X}^{\prime})$ for all $y\in\mathbb{R}$. For bivariate stochastically increasing distributions from the same Fréchet class, the conditional convex order coincides with the concordance order, as we show in Proposition 5.1. Recall that a bivariate distribution (function) is stochastically increasing (SI) if there exists a bivariate random vector $(W_{1},W_{2})$ with distribution (function) such that $W_{1}$ is SI in $W_{2}$, i.e., $\mathbb{E}[f(W_{1})\mid W_{2}=t]$ is increasing in $t$ for all increasing functions $f$ such that the expectations exist.
  2. (b)
    The conditional convex order is defined by comparing conditional survival probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order for all $v\in(0,1)$. Equivalent versions can be achieved for strict survival probabilities and cumulative probabilities. For example, $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is equivalent to
    $P(Y\leq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})\quad\text{for }\lambda\text{-\emph{almost all }}v\in(0,1);$ (4)
    see Proposition A.4. The exceptional null sets in (4) are required due to possibly different continuity properties of the transformations $F_{Y}\circ q_{Y}$ and $F_{Y^{\prime}}\circ q_{Y^{\prime}}$ at the jump points of $F_{Y}$ and $F_{Y^{\prime}}$, respectively; see Example A.5 and Proposition A.1(ii) and (iii) for details. In contrast, the transformed survival functions $t\mapsto P(Y\geq q_{Y}(t))$ and $t\mapsto P(Y^{\prime}\geq q_{Y^{\prime}}(t))$ are left-continuous and, hence, they coincide under the marginal constraint for all $t$. This motivated us to define the conditional convex order through survival probabilities.
1.3 Main results
The following main result shows that the conditional convex order in Definition 1.2 satisfies all the proposed axioms for a global dependence ordering.
Theorem 1.4 (Fundamental properties of $\preccurlyeq_{ccx}$)
For a closed set $A\subseteq[0,1]\,$, define $\mathcal{R}_{A}:=\{(Y,\mathbb{X})\colon\Omega\to\mathbb{R}^{1+p}\mid\overline{\mathsf{Ran}(F_{Y})}=A\,,\,p\in\mathbb{N}\}\,.$ Then, for $\mathcal{R}=\mathcal{R}_{A}$, the conditional convex order $\preccurlyeq_{ccx}$ in Definition 1.2 satisfies the Axioms 18.
Remark 1.5
  1. (a)
    The class $\mathcal{R}_{A}$ in Theorem 1.4 consists of all random vectors $(Y,\mathbb{X})$ that are at least two-dimensional and satisfy the marginal constraint with respect to the set $A$. For example, if $A=[0,1]$, then $\mathcal{R}_{A}$ includes all random vectors $(Y,\mathbb{X})$ such that $Y$ has a continuous distribution function; note that there is no assumption on the distribution or dimension of $\mathbb{X}$. The set $\mathcal{R}_{A}$ includes the extreme elements in conditional convex order, for example, the comonotonic random vector $(Y,Y)$ as maximal element and the independent random vector $(Y,c)$, $c\in\mathbb{R}$, as minimal element.
  2. (b)
    Under the marginal constraint, independent and perfectly dependent random vectors are globally extreme elements in $\preccurlyeq_{ccx}$ and are thus comparable to all random vectors in $\mathcal{R}_{A}$. Such a property does not hold for the regression dependence orders in [22] where independent random variables may not be comparable to other random vectors even if they have the same marginal distributions.
  3. (c)
    The conditional convex order is not symmetric, i.e., for bivariate random vectors $(Y,X)$ and $(Y^{\prime},X^{\prime})$, the relation $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$ does not imply $(X,Y)\preccurlyeq_{ccx}(X^{\prime},Y^{\prime})$. This is consistent with the asymmetric notion of perfect dependence, where $Y$ being a function of $X$ does not imply that $X$ is a function of $Y$.
  4. (d)
    As a consequence of the information monotonicity and the characterization of conditional independence, the conditional convex order satisfies the data processing inequality $(Y,\mathbb{Z})\preccurlyeq_{ccx}(Y,\mathbb{X})$ whenever $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}\,.$ It follows that $(Y,\mathbb{h}(\mathbb{X}))\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all measurable functions $\mathbb{h}$. Further, the conditional convex order is self-equitable in the sense that $(Y,\mathbb{h}(\mathbb{X}))=_{ccx}(Y,\mathbb{X})$ for all measurable functions $\mathbb{h}$ such that $Y$ and $\mathbb{X}$ are conditionally independent given $\mathbb{h}(\mathbb{X})\,.$
In the next theorem, we characterize the conditional convex order through a comparison of bivariate SI random vectors in concordance order $\leq_{c}$ (defined in (27)). To this end, consider for $(Y,\mathbb{X})$ the rearranged quantile transform $(0,1)^{2}\ni(t,u)\mapsto q_{Y|\mathbb{X}}^{\uparrow u}(t)\in\mathbb{R}$ defined in Equation (10) below. We will consider reduced random vectors of the form $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$, which are bivariate random vectors that have the same degree of dependence as $(Y,\mathbb{X})$ in conditional convex order. We refer to Proposition 2.7 for their basic properties. We denote by $U(0,1)$ the uniform distribution on $(0,1)$, assuming that the underlying probability space $(\Omega,\mathcal{A},P)$ is non-atomic and thus rich enough to ensure the existence of $U(0,1)$-distributed random variables.
Theorem 1.6 (Characterization of $\preccurlyeq_{ccx}$ by bivariate concordance order)
For $U,V\sim U(0,1)$ independent, the following statements are equivalent:
  1. (i)
    $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,$,
  2. (ii)
    $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V),U)\,.$
In particular, $(Y,\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is equivalent to $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V),U)\,.$
Remark 1.7
  1. (i)
    The concordance order requires equal marginal distributions; hence, we have
    $F_{Y}(q_{Y|\mathbb{X}}^{\uparrow U}(V))\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V))$ in Theorem 1.6(ii). This condition is equivalent to the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ resulting from the definition of the conditional convex order; to see this, combine Proposition 2.7(i) and Lemma A.2.
  2. (ii)
    If $F_{Y}$ is continuous, then $F_{Y}(q_{Y|\mathbb{X}}^{\uparrow U}(V))$ and $F_{Y^{\prime}}(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V))$ are uniform on $(0,1)$, and the distribution functions of the reduced random vectors in Theorem 1.6(ii) are SI copulas. Consequently, the conditional convex order can be reduced to a pointwise comparison of SI copulas. In particular, Theorem 1.6 strengthens Axiom 8 in the sense that the conditional convex order only depends on the bivariate SI copula obtained from the dimension reduction principle in Proposition 2.7.
The following theorem shows that the strength of functional dependence in the additive error model (1) decreases in $\sigma$ with respect to the conditional convex order. Hence, as a consequence of Theorem 4.1, Chatterjee’s rank correlation and related dependence measures are decreasing in the parameter $\sigma$ of the additive error model.
Theorem 1.8 (Additive error models)
Consider $Y=f(\mathbb{X})+\sigma\varepsilon$ and $Y^{\prime}=f(\mathbb{X})+\sigma^{\prime}\varepsilon$ for $0\leq\sigma<\sigma^{\prime}.$ Assume that $F_{Y}$ and $F_{Y^{\prime}}$ are continuous and that $\varepsilon$ is independent from $f(\mathbb{X})\,.$ Then we have $(Y,\mathbb{X})\succcurlyeq_{ccx}(Y^{\prime},\mathbb{X}).$
Remark 1.9
  1. (a)
    Besides the continuity assumption on $Y$ and the independence condition between $f(\mathbb{X})$ and $\varepsilon$, there are no assumptions on the distributions of the underlying random variables in Theorem 1.8. In particular, the error $\varepsilon$ is not assumed to follow a symmetric distribution, to be infinitely divisible, or to have zero mean. Extensions to non-additive error models are studied in Section 3.2.
  2. (b)
    The proof of Theorem 1.8 is based on Theorem 2.2 which characterizes the conditional convex order by decreasing rearrangements of transformed conditional distribution functions. If $Y$ follows a convolution semigroup, the behavior of $\xi$ can be verified much easier using the data processing inequality for $\xi$ in [17]: In this case, $Y$ admits a representation $Y=X_{1}+\varepsilon_{1}+\varepsilon_{2}$ for independent random variables $X_{1},\varepsilon_{1},\varepsilon_{2}$; see e.g. [38, Theorem 14.50]. For $X_{2}:=X_{1}+\varepsilon_{1}$, it follows that $(Y\mid X_{2}=x)\sim\varepsilon_{2}+x$ and $(X_{1}\mid X_{2}=x)\sim x-\varepsilon_{1}$. Hence, $Y$ and $X_{1}$ are conditionally independent given $X_{2}$, which implies $\xi(Y,X_{1})\leq\xi(Y,X_{2})$. More generally, the data processing inequality for the conditional convex order yields $(Y,X_{1})\preccurlyeq_{ccx}(Y,X_{2})$; see Remark 1.5(d).
1.4 Related literature and our contribution
The growing need to evaluate the significance of features in large and complex datasets has, in recent years, motivated research on association measures satisfying various desirable axioms, as studied in [13, 37, 44, 49, 53]. Dependence measures that characterize independence and perfect functional dependence have been studied by Dette et al. [22] and Gamboa et al. [26], and recently in [5, 7, 8, 13, 15, 21, 29, 61, 62]. In their seminal work, Azadkia and Chatterjee [7] developed a nearest-neighbor-based estimator for the rank correlation $\xi$ in (2), and they established a powerful, model-free variable selection method. Their method relies on the fact that $\xi$ not only satisfies the zero-independence and max-functionality property but also characterizes conditional independence, fulfills information monotonicity, and satisfies several invariance properties.
While measures of association like the Pearson correlation, Spearman’s rho or Kendall’s tau have a clear interpretation, there is still lack in interpreting dependence measures like $\xi$. Closed-form expressions are only rarely available for some specific distributions and measures; see e.g. [2] for a closed-form expression of $\xi$ for the multivariate normal distribution. To better understand the properties and behaviors of such dependence measures, an ordering that underlies these measures is of fundamental importance. It is well known that measures of concordance, such as Kendall’s $\tau$, Spearman’s $\rho$, or Gini’s $\gamma$, are increasing in the concordance order [32, Section 2.12]. In contrast, for dependence measures like Chatterjee’s $\xi$ an open problem concerns the identification of a dependence ordering that underlies these measures. We resolve this problem and contribute to the literature on dependence measures as follows:
  • We propose Axioms 18 as natural and desirable axioms for a global dependence ordering that describes the strength of functional dependence of a random variable $Y$ on a random vector $\mathbb{X}$. As we show in Theorem 1.4, the conditional convex order introduced in Definition 1.2 satisfies all these axioms. To the best of our knowledge, there is no work that addresses all these axioms, and we are not aware of any dependence ordering that has all these properties. For an overview of dependence orderings, we refer to [32, 47, 57].
  • In Section 2, we derive two characterizations of the conditional convex order. The first one relies on the Schur order used to compare the variability of conditional distribution functions; see Theorem 2.2. As an example, we verify $\preccurlyeq_{ccx}$ for bivariate Bernoulli distributions; see Proposition 2.5. For the second characterization, we introduce a dimension reduction principle to transform a random vector $(Y,\mathbb{X})$ to a bivariate SI random vector. Due to Proposition 2.7, this reduced random vector has the same strength of functional dependence in the sense of the conditional convex order. In our second main result, Theorem 1.6, we characterize $\preccurlyeq_{ccx}$ by comparing reduced random vectors in concordance order.
  • In Section 3, we provide sufficient conditions for verifying $\preccurlyeq_{ccx}$ in the additive error model (1) (Theorem 1.8) and for various copula-based models (Theorem 3.3). For the multivariate normal distribution, we obtain a notable characterization of the conditional convex order and show that multivariate normal random vectors are always comparable in $\preccurlyeq_{ccx}$ (Theorem 3.1).
  • As shown by Borgonovo et al. [13], convexity is essential for a dependence measure to satisfy information monotonicity and to characterize independence and perfect dependence. In Theorem 4.1 and Theorem 4.4, we use the conditional convex order to generate new dependence measures that are increasing in $\preccurlyeq_{ccx}$, possess the zero-independence and max-functionality property, characterize conditional independence, satisfy information monotonicity, and exhibit various useful invariance properties. We are not aware of related construction methods that yield dependence measures with all these properties.
  • In Section 5, we compare the conditional convex order with related positive and global dependence orderings from the literature [22, 58, 59, 63, 68]. We show that none of the existing orderings satisfies Axioms 18, nor do they describe the behavior of Chatterjee’s rank correlation across various models.
Throughout the paper, boldface symbols $\mathbb{x}$ and $\mathbb{X}$ are used for vectors and random vectors, respectively, and standard symbols $y$ and $Y$ for real numbers and real-valued random variables. All proofs are deferred to the Appendix B.
2 Characterizations of $\preccurlyeq_{ccx}$
In this section, we derive two characterizations of the conditional convex order. The first is based on the Schur order, which compares the variability of integrable functions via their decreasing rearrangements. The second characterization is formulated in terms of the concordance order applied to dimension-reduced bivariate SI random vectors.
2.1 Decreasing rearrangements
For verifying the conditional convex order, we use the Hardy-Littlewood-Polya theorem which characterizes the standard convex order by the Schur order for functions. The latter is defined by comparing decreasing rearrangements as follows. For $d,d^{\prime}\in\mathbb{N}\,$, let $f\colon(0,1)^{d}\to\mathbb{R}$ and $g\colon(0,1)^{d^{\prime}}\to\mathbb{R}$ be integrable functions. Then the Schur order $f\prec_{S}g$ is defined by
$\int_{0}^{x}f^{*}(t)\mathsf{\,d}t$ $\leq\int_{0}^{x}g^{*}(t)\mathsf{\,d}t\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{for all }x\in(0,1)\text{ and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \int_{0}^{1}f^{*}(t)\mathsf{\,d}t=\int_{0}^{1}g^{*}(t)\mathsf{\,d}t\,$, (5)
where $h^{*}\colon(0,1)\to\mathbb{R}$ denotes the decreasing rearrangement of an integrable function $h\colon(0,1)^{k}\to\mathbb{R}\,$, $k\in\mathbb{N}\,$, i.e., the essentially (with respect to the Lebesgue measure) uniquely determined decreasing function $h^{*}$ such that $\lambda(h^{*}\geq w)=\lambda^{k}(h\geq w)$ for all $w\in\mathbb{R}\,;$ see [19]. Here, $\lambda^{k}$ denotes the Lebesgue measure on $(0,1)^{k}$.
The Schur order is characterized by the following multivariate version of the Hardy-Littlewood-Polya theorem; see [19, Theorem 2.5].
Lemma 2.1 (Hardy-Littlewood-Polya theorem)
Let $f\colon(0,1)^{d}\to\mathbb{R}$ and $g\colon(0,1)^{d^{\prime}}\to\mathbb{R}$ be integrable functions. Then the following statements are equivalent:
  1. (a)
    $f\prec_{S}g$
  2. (b)
    $\int_{(0,1)^{d}}\varphi\circ f\mathsf{\,d}\lambda^{d}\leq\int_{(0,1)^{d^{\prime}}}\varphi\circ g\mathsf{\,d}\lambda^{d^{\prime}}$ for all convex functions $\varphi\colon\mathbb{R}\to\mathbb{R}$ such that the integrals exist.
To apply the Hardy-Littlewood-Polya theorem for a characterization $\preccurlyeq_{ccx}$, we use quantile transformations to get back to the unit interval/hypercube. Recall that $q_{Y}$ is the left-continuous quantile function of $Y$. We also need the multivariate quantile transform ${\bf q}_{\mathbb{X}}\colon(0,1)^{p}\to\mathbb{R}^{p}$ defined in (36). It satisfies $\mathbb{X}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}{\bf q}_{\mathbb{X}}(\mathbb{U})$ for $\mathbb{U}$ uniform on $(0,1)^{p}$. Now, define for $v\in(0,1)$ the function $\eta_{Y|\mathbb{X}}^{v}\colon(0,1)^{p}\to\mathbb{R}$ by
$\eta_{Y|\mathbb{X}}^{v}(\mathbb{u}):=$ $P(Y\leq q_{Y}(v)\mid\mathbb{X}={\bf q}_{\mathbb{X}}(\mathbb{u}))\,.$ (6)
Due to the following result, the conditional convex order can be characterized by a comparison of the functions $\eta_{Y|\mathbb{X}}^{v}$ and $\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v}$ in Schur order. We will use this characterization to verify the conditional convex order in various models; in particular, Theorem 1.8 on additive error models and Proposition 2.5 on bivariate Bernoulli distributions rely on it.
Theorem 2.2 (Characterization of $\preccurlyeq_{ccx}$ by the Schur order)
The following statements are equivalent:
  1. (i)
    $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$
  2. (ii)
    $\eta_{Y|\mathbb{X}}^{v}\prec_{S}\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v}$ for $\lambda$-almost all $v\in(0,1)$.
Remark 2.3
  1. (a)
    In our preliminary work [3, 5], we defined the so-called Schur order for conditional distributions by Theorem 2.2(ii) to establish ordering results for Chatterjee’s rank correlation in the setting of bivariate copula models. Here, we consider the conditional convex order, which is equivalent to the Schur order for conditional distributions but conceptually simpler.
  2. (b)
    Global dependence orderings based on the Schur order in (5) are also studied in [30, 54]. In the first reference, densities are compared in the Schur order to analyze the strength of dependence within multivariate distributions. The second reference studies so-called Lorentz curves (which are related to the Schur order) to compare likelihood ratios between probability measures with finite support. In contrast, in our approach, we compare conditional distribution functions and analyze the strength of dependence of $Y$ on $\mathbb{X}$. Note that the definition of $\preccurlyeq_{ccx}$ does not require assumptions like absolute continuity or distributions with finite support.
In the following example, we determine the decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (6) for bivariate Bernoulli distributions.
Example 2.4 (Bivariate Bernoulli distribution)
Consider a bivariate Bernoulli distributed random vector $(Y,X)$ taking on values $(0,0),(0,1),(1,0)$ and $(1,1)$. Denote by $p_{ji}:=P(Y=j,X=i)$, $j,i\in\{0,1\}$, the corresponding probability of occurrence such that $q:=P(Y=1)=p_{10}+p_{11}\in(0,1)$ and $p:=P(X=1)=p_{01}+p_{11}\in(0,1)$. With this notation, the bivariate Bernoulli distribution has in total $6$ parameters with $3$ degrees of freedom, where $q$ determines the distribution of $Y$ and where $\alpha:=p_{00}/(1-p)$ and $\beta:=p_{01}/p$ determine the conditional probabilities $P(Y=0\mid X=0)$ and $P(Y=0\mid X=1)$. The conditional distribution function $\eta_{Y|X}^{v}$ in (6) and its decreasing rearrangement $(\eta_{Y|X}^{v})^{\ast}$ in (9) are given by
$\eta_{Y|X}^{v}(u)$ $=\begin{cases}\alpha\,\mathds{1}_{(0,1-p]}(u)+\beta\,\mathds{1}_{(1-p,1]}(u)&\text{if }v\in(0,1-q],\\ 1&\text{if }v\in(1-q,1],\end{cases}$ (7)
$(\eta_{Y|X}^{v})^{*}(u)$ $=\begin{cases}(\alpha\vee\beta)\,\mathds{1}_{(0,z]}(u)+(\alpha\wedge\beta)\,\mathds{1}_{(z,1]}(u)&\text{if }v\in(0,1-q],\\ 1&\text{if }v\in(1-q,1],\end{cases}$ (8)
where z:=1pz:=1-p if $\alpha\geq\beta$ and $z:=p$ otherwise. Here, $\vee$ and $\wedge$ denote the maximum and minimum of two real numbers. The integrated decreasing rearrangement $x\mapsto\int_{0}^{x}(\eta_{Y|X}^{v})^{\ast}(u)\mathsf{\,d}u$ in (5) is piecewise linear with a breakpoint at $z$ and slope $\alpha\vee\beta$ on $[0,z]$ and slope $\alpha\wedge\beta$ on $(z,1]$; see Figure 1 for an illustration.
Figure 1: The red curve depicts the piecewise linear integrated decreasing rearrangement $x\mapsto\int_{0}^{x}(\eta_{Y|X}^{v})^{\ast}(u)\mathsf{\,d}u$, $v\in(0,1-q]$, of the Bernoulli-distributed random vector $(Y,X)$ discussed in Example 2.4. For any Bernoulli-distributed random vector $(Y^{\prime},X^{\prime})$ with $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$ or, equivalently, $\eta_{Y|X}^{v}\prec_{S}\eta_{Y^{\prime}|X^{\prime}}^{v}$ for Lebesgue almost all $v\in(0,1-q]$, its integrated decreasing rearrangement (blue curve) lies above the red curve and hence must have a slope greater than $\alpha\vee\beta$ in the interval starting at 0 and a slope less than $\alpha\wedge\beta$ in the interval ending at $1$. Due to the marginal constraint both the red and blue functions equal 1q1-q at point 11.
The following result characterizes the conditional convex order for bivariate Bernoulli-distributed random vectors.
Proposition 2.5 ($\preccurlyeq_{ccx}$-ordering of bivariate Bernoulli distributions)
Let $(Y,X)$ and $(Y^{\prime},X^{\prime})$ be Bernoulli distributed random vectors with parameters as in Example 2.4. Then the following statements are equivalent:
  1. (i)
    $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$,
  2. (ii)
    $q=q^{\prime}$ as well as $\alpha^{\prime}\wedge\beta^{\prime}\leq\alpha\wedge\beta$ and $\alpha\vee\beta\leq\alpha^{\prime}\vee\beta^{\prime}$.
Further, we have that
  1. (iii)
    $X$ and $Y$ are independent if and only if $\alpha=\beta$.
  2. (iv)
    $Y$ perfectly depends on $X$ if and only if $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$.
  3. (v)
    $X$ and $Y$ are comonotone if and only if $\alpha=\tfrac{1-q}{1-p}\wedge 1$.
  4. (vi)
    $X$ and $Y$ are countermonotone if and only if $\alpha=0\vee(1-\tfrac{q}{1-p})$.
Remark 2.6
Consider the setting in Proposition 2.5.
  1. (a)
    In (ii), the condition $q=q^{\prime}$ follows from the marginal constraint of the conditional convex order, i.e., $\{0,1-q,1\}=\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}=\{0,1-q^{\prime},1\}$. Large and small values of $\alpha$ and $\beta$ yield strong couplings between $Y$ and $X$, and hence a stronger functional dependence in conditional convex order than moderate values of $\alpha$ and $\beta$.
  2. (b)
    There are several alternatives to characterize independence, for example, $\alpha=1-q$.
  3. (c)
    In the characterization of perfect dependence in (iv), the condition $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$ corresponds to the case $Y=X$ or Y=1XY=1-X almost surely. In these cases, it is $p=q$ and p=1qp=1-q, respectively. Due to (v) and (vi), perfect dependence of $Y$ on $X$ implies comonotonicity or countermonotonicity of $(Y,X)$; however, the converse is not true because $(Y,X)$ being comonotone (resp. countermonotone) only means that $Y=q_{Y}(U)$ and $X=q_{X}(U)$ (resp. $X=q_{X}(1-U)$) a.s. for $U$ uniform on $(0,1)$. Consequently, the extreme cases of $\leq_{c}$ and $\preccurlyeq_{ccx}$ generally differ.
  4. (d)
    $Y$ is SI in $X$ if and only if $\alpha\geq\beta$. If, additionally, $p=p^{\prime}$, $q=q^{\prime}$, and $\alpha^{\prime}\geq\beta^{\prime}$, then we have $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$ $\Longleftrightarrow$ $\alpha\leq\alpha^{\prime}$ $\Longleftrightarrow$ $p_{00}\leq p_{00}^{\prime}$ $\Longleftrightarrow$ $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$. This confirms Proposition 5.1 in the case of Bernoulli distributions.
2.2 A Dimension reduction principle
In this subsection, we derive a characterization of the conditional convex order in terms of the concordance order for bivariate SI distributions. To this end, we perform a dimension reduction based on decreasing rearrangements. For convenience, we work with cumulative probabilities and distribution functions instead of survival probabilities and survival functions.
To derive various properties of the conditional convex order, we denote, for all $y\in\mathbb{R}$, by $u\mapsto F_{u}(y)$ the decreasing rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. Recall that $u\mapsto F_{u}(y)$ is $\lambda$-almost surely uniquely determined. In the sequel, we will consider versions such that $y\mapsto F_{u}(y)$ is a distribution function for all $u\in(0,1)$; see Lemma B.1 for the existence of such rearrangements. Setting $y=q_{Y}(v)$ for $v\in(0,1)$, the function $u\mapsto F_{u}(q_{Y}(v))$ is the decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (6), i.e.
$(\eta_{Y|\mathbb{X}}^{v})^{*}(u)=F_{u}(q_{Y}(v))\quad\text{for all }u\in[0,1].$ (9)
Using that $F_{u}$ is a distribution function on $\mathbb{R}$, we define its associated quantile function by
$q_{Y|\mathbb{X}}^{\uparrow u}(t):=F_{u}^{-1}(t),\quad t\in(0,1],$ (10)
which we refer to as rearranged quantile transform. To explain the choice of the notation in (10), let $V$ be uniform on $(0,1)$. First, note that $q_{Y|\mathbb{X}}^{\uparrow u}(V)$ is a random variable with distribution function $F_{u}$. Since, by construction, $F_{u}(y)$ is decreasing in $u$ for all $y\in\mathbb{R}$, the family $\{q_{Y|\mathbb{X}}^{\uparrow u}(V)\}_{u\in[0,1]}$ of random variables induced by (10) is increasing in $u$ with respect to the stochastic order. This means that $u_{1}\leq u_{2}$ implies $\mathbb{E}f(q_{Y|\mathbb{X}}^{\uparrow u_{1}}(V))\leq\mathbb{E}f(q_{Y|\mathbb{X}}^{\uparrow u_{2}}(V))$ for all increasing functions $f$ such that the expectations exist.
[…]
Figure 2: Example for the construction of the bivariate SI random vector $(q_{Y|X}^{\uparrow U}(V),U)$ in Proposition 2.7(ii) when a discrete bivariate random vector $(Y,X)$ mapping into a grid $\{a_{1},\ldots,a_{4}\}\times\{b_{1},\ldots,b_{4}\}$ is given: The top left matrix illustrates the mass distribution of $(Y,X)$; the bottom left matrix describes the associated conditional distribution functions $\eta_{Y|X}^{v}\,;$ the bottom right matrix defines the conditional distribution function $v\mapsto F_{u}(q_{Y}(v))$ that is defined in (9) and decreasing in $u$; the top right matrix describes the mass distribution of the SI random vector $(q_{Y|X}^{\uparrow U}(V),U)$ defined by (10).
Using the rearranged quantile transform constructed above, the following result shows that every random vector $(Y,\mathbb{X})$ can be reduced to a bivariate SI random vector having the same strength of functional dependence in the sense of the conditional convex order.
Proposition 2.7 (Dimension reduction to bivariate SI distributions)
Let $U,V\sim U(0,1)$ be independent. Then the rearranged quantile transform defined by (10) has the following properties:
  1. (i)
    $q_{Y|\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y\,$,
  2. (ii)
    $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ is a bivariate SI random vector,
  3. (iii)
    $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})\,.$
Remark 2.8
  1. (a)
    Due to Proposition 2.7, the rearranged quantile transform relates a $(1+p)$-dimensional random vector $(Y,\mathbb{X})$ to a bivariate SI random vector such that the dependence information in the sense of the conditional convex order remains invariant. Consequently, the strength of functional dependence is preserved in the sense that all dependence measures which are monotone in $\preccurlyeq_{ccx}$ are invariant under this transformation. We refer to Section 4 for classes of $\preccurlyeq_{ccx}$-increasing dependence measures. Our transformation extends the rearrangement of bivariate copulas in [4, Proposition 3.17] and [62, Theorem 2.3] to arbitrary marginal distributions and to a multivariate random vector $\mathbb{X}$.
  2. (b)
    For $A={\overline{\mathsf{Ran}(F_{Y})}}$, $(Y,\mathbb{X})$ and $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ are minimal elements in $\mathcal{R}_{A}$ if and only if $Y$ and $\mathbb{X}$ are independent or, equivalently, if $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$ are independent. Similarly, they are maximal elements in $\mathcal{R}_{A}$ if and only if $Y$ perfectly depends on $\mathbb{X}$, which is equivalent to $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ being perfectly dependent on $U$. Since $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ is SI in $U$, the latter property is equivalent to comonotonicity of $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$.
For a bivariate random vector with finite support, the construction of the rearranged quantile transform in (10) is explained in Figure 2. For a multivariate normal random vector, we illustrate the dimension reduction principle and the construction of the bivariate SI random vector $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ in the following example. Note that we use the basic ideas in this example to study sufficient conditions for verifying the conditional convex order in Section 3.2.
Example 2.9 (Dimension reduction for multivariate normal distribution)
Let $(Y,\mathbb{X})\sim N(\mathbf{0},\Sigma)$ be a $(1+p)$-dimensional normal random vector. In the sequel, we construct the bivariate SI random vector $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ in Proposition 2.7 based on the dimenion reduction principle in (9) and (10). Consider the decomposition of the covariance matrix into
$\Sigma=\begin{pmatrix}\sigma_{Y}^{2}&\Sigma_{\mathbb{X},Y}\\ \Sigma_{Y,\mathbb{X}}&\Sigma_{\mathbb{X}}\end{pmatrix}$. (11)
Define $S:=A\mathbb{X}$ for $A:=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}/\sqrt{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}}\,$, where $\Sigma_{\mathbb{X}}^{-}$ denotes a generalized inverse of $\Sigma_{\mathbb{X}}$ such as the Moore-Penrose inverse. Key for the dimension reduction principle and for the construction of the rearranged quantile transform $q_{Y|\mathbb{X}}^{\uparrow u}$ in (10) is the identity
$(Y\mid\mathbb{X}=\mathbb{x})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y\mid S=A\mathbb{x})\quad\text{for }P^{\mathbb{X}}-\text{almost all }\mathbb{x}\in\mathbb{R}^{p}\,;$ (12)
see [2, Proof of Proposition 2.7]. For $k:=\mathsf{rank}(\Sigma_{\mathbb{X}})\,$, let $B$ be a $p\times k$ matrix such that $\Sigma_{\mathbb{X}}=BB^{T}.$ Then the transformed conditional distribution function $\eta_{Y|\mathbb{X}}^{v}$ in (6) is given by
$\begin{split}\eta_{Y|\mathbb{X}}^{v}(\mathbb{u})&=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid\mathbb{X}=B{\bf\Phi}^{-1}(\mathbb{u}))\\ &=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=AB{\bf\Phi}^{-1}(\mathbb{u}))\\ &=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=\Phi^{-1}(u))\end{split}$ (13)
for $u=\Phi(AB{\bf\Phi}^{-1}(\mathbb{u}))\,$, where ${\bf\Phi}^{-1}(\mathbb{u})$ denotes a column vector with the standard normal quantile function applied to each component of $\mathbb{u}\in(0,1)^{k}\,.$ Note that $(Y,S)$ is bivariate normal with correlation
$\mathrm{Cor}(Y,S)=\frac{\operatorname{Cov}(Y,A\mathbb{X})}{\sqrt{\mathrm{Var}(Y)\mathrm{Var}(A\mathbb{X})}}=\sqrt{\frac{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}}{\sigma_{Y}^{2}}}\geq 0,$ (14)
which implies that $Y$ is SI in S;S\,; see e.g. [50]. Let $\mathbb{U}$ be a random vector that is uniform on $(0,1)^{k}\,$, and define $U:=\Phi(S)=\Phi(AB{\bf\Phi}^{-1}(\mathbb{U}))$. Then $U$ is uniform on $(0,1)$ because $S$ is standard normal. The decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (13) is given by
$(\eta_{Y|\mathbb{X}}^{v})^{*}(u)=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=\Phi^{-1}(u))$ (15)
since the right-hand side of (15) is decreasing in $u$ and
$\begin{split}\lambda(\{u\mid(\eta_{Y|\mathbb{X}}^{v})^{*}(u)\geq z\})&=P(\{\omega\mid(\eta_{Y|\mathbb{X}}^{v})^{*}(U(\omega))\geq z\})\\ &=P(\{\omega\mid\eta_{Y|\mathbb{X}}^{v}(\mathbb{U}(\omega))\geq z\})=\lambda^{p}(\{\mathbb{u}\mid\eta_{Y|\mathbb{X}}^{v}(\mathbb{u})\geq z\})\end{split}$ (16)
for all $z\in\mathbb{R}\,.$ The second equality in (16) follows from (13). For the first equality in (16), we use that $U$ is uniform on (0,1);(0,1)\,; for the third equality we use that $\mathbb{U}$ is uniform on $(0,1)^{k}\,.$ For fixed $u\in(0,1)\,$, the rearranged distribution function $F_{u}$ in (9) is given by
$F_{u}(y)=P(Y\leq y\mid S=\Phi^{-1}(u))=(\eta_{Y|\mathbb{X}}^{v})^{*}(u)$ (17)
for $y=q_{Y}(v)=\sigma_{Y}\Phi^{-1}(v)\,.$ Hence, the rearranged quantile transform defined in (10) is
$q_{Y|\mathbb{X}}^{\uparrow u}(t)=F_{u}^{-1}(t)=F_{Y|S=\Phi^{-1}(u)}^{-1}(t)\,,\quad t\in(0,1)\,.$ (18)
Finally, the bivariate SI random vector in Proposition 2.7(ii) is given by
$(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=(F_{Y|S=\Phi^{-1}(U)}^{-1}(V),U).$ (19)
Note that the left-hand side in (19) satisfies $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})$, which follows from (13) and (17) and verifies Proposition 2.7(iii). Due to (12), we know that $(Y,\mathbb{X})=_{ccx}(Y,S)$. Hence, the reduced random vectors in Theorem 1.6 satisfy
$(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y}\circ Y,\Phi(S))$ (20)
in the case of the multivariate normal distribution.
3 Verifying $\preccurlyeq_{ccx}$
We use the characterizations from the previous section to verify the conditional convex order in various settings. Theorem 1.8 on additive error models is an application of the Schur ordering criterion in Theorem 2.2. In Section 3.1, we use Theorem 1.6 to characterize the conditional convex order for the multivariate normal distribution. In the second part of this section, we apply the dimension reduction principle from Section 2.2 to more general dependence structures.
3.1 Multivariate normal distribution
The following result characterizes the conditional convex order in the multivariate normal model. Its proof is based on the dimension reduction principle in Example 2.9.
Theorem 3.1 (Conditional convex order for multivariate normal distribution)
Let $(Y,\mathbb{X})\sim N(\bm{\mu},\Sigma)$ and $(Y^{\prime},\mathbb{X}^{\prime})\sim N(\bm{\mu}^{\prime},\Sigma^{\prime})$ be multivariate normal random vectors with non-degenerate components $Y$ and $Y^{\prime}.$ For $\Sigma$ and $\Sigma^{\prime}$ decomposed as in (11), we have
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\Longleftrightarrow\quad\,\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}\,\leq\,\Sigma_{Y^{\prime},\mathbb{X}^{\prime}}\Sigma_{\mathbb{X}^{\prime}}^{-}\Sigma_{\mathbb{X}^{\prime},Y^{\prime}}/\sigma_{Y^{\prime}}^{2}.$ (21)
Remark 3.2
  1. (a)
    Due to Theorem 3.1, multivariate normal random vectors $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$ are always comparable in conditional convex order, provided $\sigma_{Y}$ and $\sigma_{Y^{\prime}}$ are positive. It is neither assumed that the random vectors have the same dimension nor that the covariance matrices are positive definite. No conditions are imposed on $\bm{\mu}$ and $\bm{\mu}^{\prime}$ either. Due to [2, Proposition 2.7], the conditions in (21) are equivalent to $\xi(Y,\mathbb{X})\leq\xi(Y^{\prime},\mathbb{X}^{\prime})$.
  2. (b)
    In Theorem 3.1, the conditional convex order reduces to comparing the Schur complements of the covariance matrices, if $\sigma_{Y}=\sigma_{Y^{\prime}}$. Since $\mathbb{E}[Y\mid\mathbb{X}=\mathbb{x}]=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\mathbb{x}$, the conditional convex order in (21) also underlies the fraction of explained variance $\mathrm{Var}(\mathbb{E}[Y\mid\mathbb{X}])/\mathrm{Var}(Y)=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}$ as well as the kernel partial correlation for a linear kernel in [29, Remark 18].
In Example 3.5 below, we verify the ordering condition in (21) for the equicorrelated normal distribution and the multivariate normal distribution with independent predictor variables.
3.2 Sufficient dimension reduction
To derive sufficient conditions for verifying $\preccurlyeq_{ccx}$, we generalize the dimension reduction principle discussed for the multivariate normal distribution in Example 2.9. To this end, we extend the identity in (12) to general transformations and denote a function $g\colon\mathbb{R}^{p}\to\mathbb{R}$ as sufficient for (the conditional distribution) $Y|\mathbb{X}$ if
$(Y\mid\mathbb{X}=\mathbb{x})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y\mid g(\mathbb{X})=g(\mathbb{x}))\quad\text{for }P^{\mathbb{X}}\text{-almost all }\mathbb{x}\in\mathbb{R}^{p}\,.$ (22)
The terminology of $g$ being sufficient for $Y|\mathbb{X}$ coincides with the classical concept of a sufficient statistic $g$ for the family of conditional distributions $\{P^{Y|\mathbb{X}=\mathbb{x}}\mid\mathbb{x}\in\mathbb{R}^{p}\}$. For a linear function $g,$ the concept in (22) is also referred to as sufficient for dimension reduction in [18, 41, 42, 40]. An example of a sufficient function for $Y|\mathbb{X}$ is the function $f$ in the additive error model (1) or the linear transformation $f(\mathbb{x})=A\mathbb{x}$ in Example 2.9. Note that (22) is equivalent to conditional independence of $Y$ and $\mathbb{X}$ given $g(\mathbb{X})$; see e.g. [33, Proposition 6.6].
In the following theorem, we provide simple conditions for verifying the conditional convex order under some stochastic monotonicity assumptions. Recall that we write $Y\uparrow_{st}Z$ if $Y$ is SI in $Z$, and similarly, $Y\downarrow_{st}Z$ if $Y$ is stochastically decreasing in $Z$ (i.e., $Y\uparrow_{st}Z$).
Theorem 3.3 ($\preccurlyeq_{ccx}$-order for copula-based models)
Assume that $g\colon\mathbb{R}^{p}\to\mathbb{R}$ is sufficient for $Y\mid\mathbb{X}$ and that $h\colon\mathbb{R}^{p^{\prime}}\to\mathbb{R}$ is sufficient for $Y^{\prime}\mid\mathbb{X}^{\prime}\,.$ Then the following statements hold true:
  1. (i)
    If $(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))\,$, then $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$
  2. (ii)
    Assume that $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}\,$, $Y\uparrow_{st}g(\mathbb{X})$, and $Y^{\prime}\uparrow_{st}h(\mathbb{X}^{\prime})\,.$ Then $C_{Y,g(\mathbb{X})}\leq_{c}C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ implies $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$
  3. (iii)
    Assume that $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}\,$, $Y\downarrow_{st}g(\mathbb{X})$, and $Y^{\prime}\downarrow_{st}h(\mathbb{X}^{\prime})\,.$ Then $C_{Y,g(\mathbb{X})}\leq_{c}C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ implies $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$
Remark 3.4
The condition $Y\uparrow_{st}g(\mathbb{X})$ is satisfied if there exists a representation $Y=f(g(\mathbb{X}),\varepsilon)$ a.s. for some componentwise increasing function $f$ and an error $\varepsilon$ that is independent of $g(\mathbb{X})$; see e.g. [45, Lemma 3.1]. Without loss of generality, $\varepsilon$ may be uniform on $(0,1)$ or standard normal. The additive error model (1) admits such a representation. Note that the proof of Theorem 1.8 applies Theorem 3.3(i) but not the criteria on pointwise comparison of copulas in parts (ii) and (iii) because the distributions of $g(\mathbb{X})$ and $\varepsilon$ in the additive error model are not given and thus the copula $C_{Y,g(\mathbb{X})}$ is not known.
In the next example, we use Theorem 3.3 to verify the $\preccurlyeq_{ccx}$-ordering criterion in (21) for two specific families of multivariate normal distributions.
Example 3.5 (Multivariate normal distribution)
Assume that $(Y,\mathbb{X})\sim N(\mathbf{0},\Sigma)$ follows a $(1+p)$-dimensional normal distribution.
  1. (a)
    Consider the case where the covariance matrix $\Sigma=(\sigma_{ij})$ is equicorrelated with $\sigma_{ij}=1$ for $i=j$ and $\sigma_{ij}=\varrho$ else for some $\varrho\in[-1/p,1]\,.$ Standard calculations show that
    $(Y\mid\mathbb{X}=\mathbb{x})\sim N\Big(\frac{\varrho}{1+(p-1)\varrho}\sum x_{i}\,,\penalty 10000\ 1-\frac{p\varrho^{2}}{1+(p-1)\varrho}\Big)\,.$ (23)
    The conditional distribution in (23) depends only on $g(\mathbb{x})=\sum x_{i}\,$, so $g$ is sufficient for $Y|\mathbb{X}\,.$ The random vector $(Y,\sum X_{i})$ follows a bivariate normal distribution with zero mean and covariance matrix
    $\Sigma=\begin{pmatrix}1&p\varrho\\ p\varrho&p(1+(p-1)\varrho)\end{pmatrix}\,$.
    The copula of $(Y,\sum X_{i})$ is the Gaussian copula with correlation $\varrho\,.$ Since the Gaussian copula family is $\leq_{c}$-increasing in its parameter and since $Y\uparrow_{st}\sum X_{i}$ for $\varrho\geq 0$, we obtain from Theorem 3.3(ii) that $(Y,\mathbb{X})$ is increasing in $\varrho$ with respect to the conditional convex order. Using radial symmetry of the bivariate normal distribution (i.e., $(-Y,-\sum X_{i})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y,\sum X_{i})$), using $Y\downarrow_{st}\sum X_{i}$ for $\varrho\leq 0$, and applying Theorem 3.3(iii), we obtain that $(Y,\mathbb{X})$ is increasing eventually in $|\varrho|$ with respect to $\preccurlyeq_{ccx}$. Note that, by (21), this means that $\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}$ is increasing in $|\varrho|$.
  2. (b)
    Consider the case where the components of $\mathbb{X}=(X_{1},\ldots,X_{p})$ are independent and where $\mathrm{Cor}(Y,X_{i})=\varrho$ for all $i$. Then the covariance matrix $\Sigma$ is positive semi-definite if and only if $\varrho\in[-p^{-1/2},p^{-1/2}]$. Again, the conditional distribution
    $(Y\mid\mathbb{X}=\mathbb{x})\sim N\big(\varrho\sum x_{i}\,,\penalty 10000\ 1-p\varrho^{2}\big)\,$, (24)
    depends only on $g(\mathbb{x})=\sum x_{i}\,.$ Straightforward calculations show that
    $\big(Y,\sum X_{i}\big)\sim N\big(0,\left(\begin{smallmatrix}1&p\varrho\\ p\varrho&p\end{smallmatrix}\right)\big)\,.$
    Similar as above, we obtain from Theorem 3.3(ii),(iii) that $(Y,\mathbb{X})$ is increasing in $|\varrho|$ with respect to the conditional convex order.
In the following example, we outline various well known SI copula families that are increasing in the concordance order. This yields families of models $Y=f(g(\mathbb{X}),\varepsilon)$ that fulfill the assumptions of Theorem 3.3(ii) and thus are comparable in conditional convex order; see Remark 3.4. By some symmetry arguments, similar results can be obtained for stochastically decreasing copulas.
Example 3.6 (Copula-based models)
Let $Y=f(g(\mathbb{X}),\varepsilon)$ with copula $C_{1}=C_{Y,g(\mathbb{X})}$ and $Y^{\prime}=f(g(\mathbb{X}^{\prime}),\varepsilon^{\prime})$ with copula $C_{2}=C_{Y^{\prime},g(\mathbb{X}^{\prime})}$. Assume that $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. Then, we can verify $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ by Theorem 3.3(ii) for Archimedean, extreme-value, and elliptical copulas as follows.
  1. (a)
    Assume that $C_{i}$ is an Archimedean copula, i.e., $C_{i}$ admits a representation $C_{i}(u,v)=\psi_{i}(\varphi_{i}(u),\varphi_{i}(v))$ for some decreasing convex function $\varphi_{i}\colon[0,1]\to[0,\infty)$ with $\varphi(1)=0$, where $\psi_{i}=\varphi_{i}^{[-1]}$ denotes the (pseudo-)inverse of $\varphi_{i}$. If $\psi_{i}^{\prime}$ is log-convex, then $C_{i}$ is SI ([46, Theorem 2.8]; further, if $\varphi_{1}\circ\psi_{2}$ is subadditive, then $C_{1}\leq_{c}C_{2}$ ([48, Theorem 4.4.2]). Hence, under these conditions, we have $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. Various Archimedean copula families that meet these conditions are given in [3, Table 3].
  2. (b)
    Assume that $C_{i}$ is an extreme-value copula, i.e., $C_{i}$ admits a representation
    $C_{i}(u,v)=(\exp(\log(uv))A_{i}(\log(v)/(\log(u)+\log(v))$ where the Pickands dependence function $A_{i}$ is convex and satisfies the constraints $\max\{t,1-t\}\leq A_{i}(t)\leq 1$ for all $t\in[0,1]$. Extreme-value copulas are always SI; moreover, $A_{1}(t)\geq A_{2}(t)$ for all $t\in(0,1)$ implies $C_{1}\leq_{c}C_{2}$; see [3, Theorem 3.4]. Various well known extreme-value copula families that meet the conditions are given in [3, Table 4].
  3. (c)
    Assume that $C_{i}$ is an elliptical copula, i.e., it is the copula (implicitly obtained by Sklar’s theorem) of an elliptically distributed bivariate random vector $\mathbb{Z}_{i}$, that is, $\mathbb{Z}_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}R_{i}A_{i}\mathbb{U}$ for a non-negative random variable $R_{i}$ that is independent of the bivariate random vector $\mathbb{U}$ which is uniformly distributed on the $2$-sphere. Here, $A_{i}$ is a $2\times 2$-matrix. Assume that $\Sigma_{i}:=A_{i}A_{i}^{T}=\begin{pmatrix}1&\rho_{i}\\ \rho_{i}&1\end{pmatrix}$ for $\rho_{i}\in[-1,1]$. Sufficient conditions on $R_{i}$ and $\rho_{i}$ for $C_{i}$ being SI are given in [1, Proposition 1.2]. Further, if $\rho_{1}\leq\rho_{2}$ then $C_{1}\leq_{c}C_{2}$. Elliptical SI copula families that meet these conditions are given in [3, Table 5].
4 Dependence measures generated by $\preccurlyeq_{ccx}$
In this section, we use the conditional convex order to generate new dependence measures. While convexity guarantees that information monotonicity is inherited from $\leq_{ccx}$ to the corresponding dependence measure, the zero-independence and max-functionality properties require additional regularity conditions; compare [5, 13]. We further provide simple conditions under which these dependence measures characterize conditional independence.
As a first class of dependence measures, we consider functionals of the form
$\xi_{\varphi}(Y,\mathbb{X}):$ $=\alpha_{\varphi}^{-1}\int_{\mathbb{R}^{p}}\int_{\mathbb{R}}\varphi(F_{Y|\mathbb{X}=\mathbb{x}}(y)-F_{Y}(y))\mathsf{\,d}P^{Y}(y)\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})\quad\text{and}$
$\Lambda_{\varphi}(Y,\mathbb{X}):$ $=\beta_{\varphi}^{-1}\int_{\mathbb{R}^{2p}}\int_{\mathbb{R}}\varphi(F_{Y|\mathbb{X}=\mathbb{x}_{1}}(y)-F_{Y|\mathbb{X}=\mathbb{x}_{2}}(y))\mathsf{\,d}P^{Y}(y)\mathsf{\,d}(P^{\mathbb{X}}\otimes P^{\mathbb{X}})(\mathbb{x}_{1},\mathbb{x}_{2})$
with normalizing constants $\alpha_{\varphi}:=\int_{\mathbb{R}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-F_{Y}(y)\right)\mathsf{\,d}P^{Y}(y)\mathsf{\,d}P^{Y}(y_{1})$ and
$\beta_{\varphi}:=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-\mathds{1}_{\{y_{2}\leq y\}}\right)\mathsf{\,d}P^{Y}(y)\mathsf{\,d}(P^{Y}\otimes P^{Y})(y_{1},y_{2})\,.$ While $\xi_{\varphi}$ compares conditional distribution functions with unconditional ones, $\Lambda_{\varphi}$ measures the sensitivity of conditional distribution functions in the conditioning variable.
By the following theorem, which is based on the Hardy–Littlewood–Polya theorem (see Lemma 2.1), convexity conditions on $\varphi$ imply that $\xi_{\varphi}$ and $\Lambda_{\varphi}$ are dependence measure that are $\preccurlyeq_{ccx}$-increasing and characterize (conditional) independence and perfect dependence. Note that, for $\varphi(x)=x^{2}$, the functionals $\xi_{\varphi}$ and $\Lambda_{\varphi}$ reduce to Chatterjee’s rank correlation $\xi$ defined in Equation (2).
Theorem 4.1 (Dependence measures generated by convexity)
Assume that $\varphi:[-1,1]\rightarrow\mathbb{R}$ is convex and strictly convex in 0 with $\varphi(0)=0$. Assume that $Y$ is non-degenerate. Then $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies
$\xi_{\varphi}(Y,\mathbb{X})\leq\xi_{\varphi}(Y^{\prime},\mathbb{X}^{\prime})\qquad\textrm{ and }\qquad\Lambda_{\varphi}(Y,\mathbb{X})\leq\Lambda_{\varphi}(Y^{\prime},\mathbb{X}^{\prime})\,.$ (25)
Further, the following statements hold true:
  1. (i)
    $\xi_{\varphi}$ and $\Lambda_{\varphi}$ attain values only in $[0,1]$.
  2. (ii)
    $\xi_{\varphi}$ and $\Lambda_{\varphi}$ have the zero-independence and max-functionality property.
  3. (iii)
    If $\varphi$ is strictly convex on [1,1][-1,1], then $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$.
Remark 4.2
Under the assumptions of Theorem 4.1,
  1. (a)
    The normalizing constants $\alpha_{\varphi}$ and $\beta_{\varphi}$ are positive. Note that the marginal constraint implies that also $Y^{\prime}$ is non-degenerate.
  2. (b)
    $\xi_{\varphi}$ and $\Lambda_{\varphi}$ are dependence measures that inherit information monotonicity and invariance properties from the conditional convex order, i.e., $\xi_{\varphi}(Y,\mathbb{X})\leq\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$, $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(g(Y),\mathbf{h}(\mathbb{X}))$ for all strictly increasing functions $g$ and bijective functions $\mathbf{h}$, and $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(F_{Y}(Y),\mathbf{F}_{\mathbb{X}}(\mathbb{X}))$, similarly for $\Lambda_{\varphi}$;
  3. (c)
    Strict convexity of $\varphi$ in $\varphi(0)=0$ is sufficient for the values 0 and $1$ to characterize independence and perfect dependence, respectively. However, strict convexity in 0 is not sufficient for $\xi_{\varphi}$ to characterize conditional independence. As a counterexample, consider $\varphi\colon t\mapsto|t|$ which is convex and strictly convex in $\varphi(0)=0$. In this case, $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ does not imply conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$; see [27, Example 5.10].
Remark 4.3
  1. (a)
    The zero-independence and max-functionality property as well as information monotonicity for $\xi_{\varphi}$ can be traced back to convexity properties for measures of statistical association as studied in [13]. In Theorem 4.1, we also provide a characterization of conditional independence and show that $\xi_{\varphi}$ is increasing in $\preccurlyeq_{ccx}$.
  2. (b)
    A variant of Chatterjee’s rank correlation is the integrated $R^{2}$ in [8], defined by
    $\nu(Y,\mathbb{X}):=\int_{\mathbb{R}}\frac{\mathrm{Var}(P(Y\leq y\,|\,\mathbb{X}))}{\mathrm{Var}(\mathds{1}_{\{Y\leq y\}})}\;\mathrm{d}\tilde{\mu}(y)\quad\textrm{for }\tilde{\mu}(B):=\frac{\mu(B\cap\tilde{S})}{\mu(\tilde{S})}\,$,
    where $\mu$ is the law of $Y$ having support $S$, and where $\tilde{S}:=S\setminus\{s_{\max}\}$ if $S$ attains a maximum $s_{\max}$, and $\tilde{S}:=S$ otherwise. Note that, in comparison to $\xi$ in (2), the integral is outside the fraction. Using the representation
    $\nu(Y,\mathbb{X})=\int_{q_{Y}^{-1}(\tilde{S})}\frac{\mathrm{Var}(P(Y\leq q_{Y}(v)\,|\,\mathbb{X}))}{\mathrm{Var}(\mathds{1}_{\{Y\leq q_{Y}(v)\}})\,(1-P^{Y}(\{s_{\rm max}\}))}\;\mathrm{d}\lambda(v)$
    and the version of $\preccurlyeq_{ccx}$ for cumulative probabilities (see Remark 1.3(b)), it follows that also $\nu$ is increasing in the conditional convex order.
For a second class of dependence measures generated by the conditional convex order, we use the characterization of $\preccurlyeq_{ccx}$ via the bivariate concordance order in Theorem 1.6(ii).
Therefore, we denote by $\mathcal{R}^{\uparrow}:=\{(T,S)\mid T,S\sim U(0,1),T\uparrow_{st}S\}$ the class of bivariate SI random vectors with marginals that are uniform on $(0,1)$. Let $\mu$ be a functional on $\mathcal{R}^{\uparrow}$ that is normalized and increasing in the concordance order in the sense that $\mu(T,S)=0$ if and only if $T$ and $S$ are independent, $\mu(T,S)=1$ if and only if $T$ and $S$ are comonotone, and $(T,S)\leq_{c}(T^{\prime},S^{\prime})$ implies $\mu(T,S)\leq\mu(T^{\prime},S^{\prime})$. We say that $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$ if $(T,S)\leq_{c}(T^{\prime},S^{\prime})$ with $(T,S)\not\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(T^{\prime},S^{\prime})$ implies $\mu(T,S)<\mu(T^{\prime},S^{\prime})$. For a multivariate random vector $(Y,\mathbb{X})\in\mathcal{R}_{[0,1]}$ and for $V,U$ i.i.d. uniform on $(0,1)$, we define
$\mu(Y,\mathbb{X}):=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U),$ (26)
where $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ is a reduced random vector associated with $(Y,\mathbb{X})$. Since $(Y,\mathbb{X})\in\mathcal{R}_{[0,1]}$, we have that $F_{Y}$ is continuous. Hence, the reduced random vector $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ has uniform marginals. Since it is also SI, it is in $\mathcal{R}^{\uparrow}$, and $\mu(Y,\mathbb{X})$ in (26) is well-defined. Recall the identity $(Y,\mathbb{X})=_{ccx}(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ due to Proposition 2.7(iii) and the distributional invariance of the conditional convex order. Consequently, the functional $\mu$ in (26) depends only on the strength of functional dependence of $Y$ on $\mathbb{X}$ in the sense of the conditional convex order. Equivalently, by Theorem 1.6, $\mu$ in (26) depends only on the strength of positive dependence of $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ in the sense of the concordance order.
Theorem 4.4 (Dependence measures generated by monotone rearrangements)
The functional $\mu$ in (26) defines a dependence measure with the following properties:
  1. (i)
    $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies $\mu(Y,\mathbb{X})\leq\mu(Y^{\prime},\mathbb{X}^{\prime})$,
  2. (ii)
    $\mu(Y,\mathbb{X})=0$ if and only if $\mathbb{X}$ and $Y$ are independent,
  3. (iii)
    $\mu(Y,\mathbb{X})=1$ if and only if $Y$ perfectly depends on $\mathbb{X}$,
  4. (iv)
    $\mu(Y,\mathbb{X})\leq\mu(Y,(\mathbb{X},\mathbb{Z}))$,
  5. (v)
    If, additionally, $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$, then $\mu(Y,\mathbb{X})=\mu(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$.
Remark 4.5
Our approach on rearranged dependence measures in Theorem 4.4 is studied in [62] for bivariate random vectors with continuous marginal distribution functions. Various examples of measures of concordance that are increasing on $\mathcal{C}^{\uparrow}$, such as Kendall’s $\tau$ and Gini’s $\gamma$, or even strictly increasing, such as Spearman’s $\rho$ or the Schweizer-Wolff measure, are given in [62, Section 2.2].
5 Comparing $\preccurlyeq_{ccx}$ with related dependence orderings
In this section, we compare the conditional convex order with known dependence orderings from the literature. First, we study the fundamentally different concepts of the concordance order and the conditional convex order; see Figure 3 and Table 1. As we show in Proposition 5.1, these orders coincide for stochastically increasing distributions. We then discuss related orderings of functional dependence studied in [22, 59] and demonstrate that the conditional convex order provides a more natural and powerful framework for ordering strengths of dependencies.
5.1 Comparison of $\preccurlyeq_{ccx}$ with the concordance order
The certainly most popular positive dependence order is the concordance order. For bivariate random vectors $(Y,X)$ and $(Y^{\prime},X^{\prime}),$ it is defined by
$(Y,X)\leq_{c}(Y^{\prime},X^{\prime})\quad:\,\Longleftrightarrow\quad(Y,X)\leq_{lo}(Y^{\prime},X^{\prime})\quad\text{and}\quad(Y,X)\leq_{uo}(Y^{\prime},X^{\prime}),$ (27)
where the lower orthant and upper orthant order on the right-hand side of (27) are defined by $P(Y\leq y,X\leq x)\leq P(Y^{\prime}\leq y,X^{\prime}\leq x)$ and $P(Y>y,X>x)\leq P(Y^{\prime}>y,X^{\prime}>x)$ for all $x,y\in\mathbb{R}$, respectively, see e.g. [32, Definition 2.4]. As a direct consequence of the definition, $(Y^{\prime},X^{\prime})\leq_{c}(Y^{\prime},X^{\prime})$ implies $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$. Note that for equal marginal distributions, the bivariate concordance order is equivalent to the lower (and also to the upper) orthant order; see Proposition B.2.
To relate the concordance order and the conditional convex order, we consider, for simplicity, bivariate random vectors $(V,U)$ and $(V^{\prime},U^{\prime})$ with marginals that are uniform on $(0,1).$ This corresponds to the setting of bivariate copulas. Using the notation $\eta_{V|U}^{v}(t)=P(V\leq v\mid U=t)$ in (6) and disintegration, the concordance order can be characterized by
$(V,U)\leq_{c}(V^{\prime},U^{\prime})\quad\Longleftrightarrow\quad\int_{0}^{u}\eta_{V|U}^{v}(t)\mathsf{\,d}t\leq\int_{0}^{u}\eta_{V^{\prime}|U^{\prime}}^{v}(t)\mathsf{\,d}t\quad\text{for all }u,v\in[0,1].$ (28)
The conditional convex order has a similar representation due to its characterization via the Schur order in Theorem 2.2. The integrand now consists of the decreasing rearrangements of $\eta_{V|U}^{v}$ and $\eta_{V^{\prime}|U^{\prime}}^{v}\,$, respectively, i.e.,
$(V,U)\preccurlyeq_{ccx}(V^{\prime},U^{\prime})\quad\Longleftrightarrow\quad\int_{0}^{u}(\eta_{V|U}^{v})^{*}(t)\mathsf{\,d}t\leq\int_{0}^{u}(\eta_{V^{\prime}|U^{\prime}}^{v})^{*}(t)\mathsf{\,d}t\quad\text{for all }u,v\in[0,1].$
The integrated decreasing rearrangements above coincide with the integrals in (28) if and only if the functions $\eta_{V|U}^{v}$ and $\eta_{V^{\prime}|U^{\prime}}^{v}$ are decreasing for all v.v\,. The latter is equivalent to the property that $V$ is SI in $U$, and $V^{\prime}$ is SI in $U^{\prime}$.
An extension of the above reasoning to arbitrary marginals yields the following result. It states that the conditional convex order and the concordance order coincide for bivariate SI random vectors from the same Fréchet class.
[…]
Figure 3: Visualization of the concordance order (left plot) and the conditional convex order (right plot) in the case of bivariate copulas. Here, $W$ and $M$ denote the lower and upper Fréchet copula and $\Pi$ the independence copula. While $W$ and $M$ are the uniquely determined minimal and maximal elements in concordance order, the conditional convex order has the independence copula as global minimal element and all perfectly dependent copulas as global maximal elements.
Proposition 5.1 (Characterization of $\preccurlyeq_{ccx}$ for bivariate SI random vectors)
Let $(Y,X)$ and $(Y^{\prime},X^{\prime})$ be bivariate SI random vectors with $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}\,.$ Then the conditional convex order and the concordance order are equivalent, i.e.,
$(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})\quad\Longleftrightarrow\quad(Y,X)\leq_{c}(Y^{\prime},X^{\prime})\,.$ (29)
In particular, $(Y,X)=_{ccx}(Y^{\prime},X^{\prime})$ is equivalent to $(Y,X)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y^{\prime},X^{\prime})\,.$
Dependence orders $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$ $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$
Domain Fréchet class $\mathcal{R}_{A}=\{(Y,\mathbb{X})\mid\overline{\mathsf{Ran}(F_{Y})}=A\}$
Transitivity
Reflexivity
Antisymmetry
Symmetry
Lower bound Countermonotonicity Independence
Upper bound Comonotonicity Perfect dependence
Invariance under increasing transformations $(a(Y),X)\leq_{c}(a(Y^{\prime}),X^{\prime}))$ $(Y,\mathbb{X})\preccurlyeq_{ccx}(a(Y^{\prime}),\mathbb{X}^{\prime})$
Invariance under decreasing transformations $(b(Y),X)\leq_{c}(b(Y^{\prime}),X^{\prime}))$ $(b(Y),\mathbb{X})\preccurlyeq_{ccx}(b(Y^{\prime}),\mathbb{X}^{\prime})$
Invariance under bijective transformations $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))$
Stability under convergence weak convergence conditional weak convergence
Table 1: Overview of basic properties of the bivariate concordance order and the conditional convex order; AA is a closed subset of $[0,1]$, $a$ and $b$ are strictly increasing/decreasing functions, $h$ is a bijective function
5.2 Comparison of $\preccurlyeq_{ccx}$ with the dilation order and dispersive order
Recall that the conditional convex order is defined through the variability of conditional survival functions in the conditioning variable. As discussed at the beginning of Section 1.2, another approach of constructing a dependence order for $\xi$ is to study the variability of the conditional survival function in the conditioned variable, i.e., to study the variability of the random variable
$P(Y\geq q_{Y}(W)\mid\mathbb{X}={\bf q}_{\mathbb{X}}(\mathbb{u}))$ (30)
for Lebesgue-almost all $\mathbb{u}\in[0,1]^{p}.$ Here, $W$ is uniform on $(0,1)$ and independent from $(Y,\mathbb{X})\,.$ Similar to the maximum property of the conditional convex order, maximal variability in (30) (in convex order) corresponds to perfect dependence of $Y$ on $\mathbb{X}$; see [22, Proposition 3(ii)]. However, a drawback of this approach is that the random variable in (30) is not constant for independent $\mathbb{X}$ and $Y$. Hence, independence of $\mathbb{X}$ and $Y$ cannot be identified with minimal elements in convex order. Moreover, a comparison of the random variables in (30) with respect to the convex order requires equal means, which is typically not the case for conditional distributions. To overcome the latter shortcoming, the idea in [22] is to center the conditional distributions by subtracting conditional means. This leads to the concept of dilation order defined by
$S$ $\leq_{dil}T\quad\text{if}\quad S-\mathbb{E}S\leq_{cx}T-\mathbb{E}T\,.$ (31)
A stronger order, that implies the dilation order, is the dispersive order, also discussed in [22]. The following result can be extended to multivariate vectors $X$ and $X^{\prime}$ with general distribution functions.
Proposition 5.2 (Dilation ordering critertion [22, Theorem 2])
Let $(Y,X)$ and $(Y^{\prime},X^{\prime})$ be bivariate random vectors with continuous marginal distribution functions. Then
$(F_{Y}(Y)|X=q_{X}(u))\leq_{dil}(F_{Y^{\prime}}(Y^{\prime})|X^{\prime}=q_{X^{\prime}}(u))\quad\text{for }\lambda\text{-almost all }u\in(0,1)$ (32)
implies $\xi(Y,X)\leq\xi(Y^{\prime},X^{\prime})\,.$
To verify the dilation order in (32), one needs to substract the conditional expectations $\mathbb{E}[F_{Y}(Y)|X=q_{X}(u)]$ and $\mathbb{E}[F_{Y^{\prime}}(Y^{\prime})|X^{\prime}=q_{X^{\prime}}(u)],$ respectively. However, as we discuss in the sequel, the centered conditional distributions are even in the simplest models not comparable in convex order, so that the dilation order in (32) cannot be verified; see Figure 4. To this end, we make use of the following necessary condition for convex order; see [57, Equation (3.A.12)]. We denote by $\mathop{\mathrm{supp}}(S)$ the support of a random variable $S$. Similarly, $\mathop{\mathrm{supp}}(S\mid T=t)$ is the support of the conditional distribution of $S$ given $T=t$.
Lemma 5.3
Let $S$ and $T$ be random variables whose supports are intervals. Then $S\leq_{cx}T$ implies $\mathop{\mathrm{supp}}(S)\subseteq\mathop{\mathrm{supp}}(T)\,.$
As a consequence of the above lemma, $S$ and $T$ are not comparable in convex order if
$\inf(\mathop{\mathrm{supp}}(S))<\inf(\mathop{\mathrm{supp}}(T))\quad\text{and}\quad\sup(\mathop{\mathrm{supp}}(S))<\sup(\mathop{\mathrm{supp}}(T))\,.$ (33)
Using (33), we obtain the following result, which limits the use of the dilation order criterion for conditional distributions in Proposition 5.2.
Proposition 5.4
Let $(V,U)$ and $(V^{\prime},U^{\prime})$ be bivariate random vectors. Assume that $\mathbb{E}[V\mid U=u]\neq\mathbb{E}[V^{\prime}\mid U^{\prime}=u]$ and $\mathop{\mathrm{supp}}(V\mid U=u)=\mathop{\mathrm{supp}}(V^{\prime}\mid U^{\prime}=u)\,.$ Then, neither $(V\mid U=u)\leq_{dil}(V^{\prime}\mid U^{\prime}=u)$ nor $(V\mid U=u)\geq_{dil}(V^{\prime}\mid U^{\prime}=u)\,.$
As studied in Section 3, the conditional convex order can be verified in various settings, including the additive error model (1). In contrast, the approach via the dilation order criterion in Proposition 5.2 is too strong even in standard settings, as the following example shows.
Example 5.5 (Dilation order)
Consider the additive error models $Y=X+\sigma\varepsilon$ and $Y^{\prime}=X+\sigma^{\prime}\varepsilon$ for $0<\sigma<\sigma^{\prime}$ with $\varepsilon,X$ standard normal and independent. To verify (32), we obtain for $W\sim U(0,1)$ that
$F_{F_{Y}(Y)|X=q_{X}(u)}^{-1}(W)$ $=\Phi\left(\frac{(1+\sigma)\Phi^{-1}(W)-\Phi^{-1}(u)}{\sigma}\right)\quad\text{and}$ (34)
$F_{F_{Y^{\prime}}(Y^{\prime})|X=q_{X}(u)}^{-1}(W)$ $=\Phi\left(\frac{(1+\sigma^{\prime})\Phi^{-1}(W)-\Phi^{-1}(u)}{\sigma^{\prime}}\right)$ (35)
for $\lambda$-almost all $u\in(0,1)\,.$ It is not difficult to see that $\mathbb{E}[F_{Y}(Y)\mid X=q_{X}(u)]\neq\mathbb{E}[F_{Y^{\prime}}(Y^{\prime})\mid X=q_{X}(u)]$ for $\lambda$-almost all $u\in(0,1)\,.$ Since $\mathop{\mathrm{supp}}(F_{Y}(Y)\mid X=q_{X}(u))=\mathop{\mathrm{supp}}(F_{Y^{\prime}}(Y^{\prime})\mid X=q_{X}(u))=[0,1]$ for all $u\in(0,1)\,$, we obtain from Proposition 5.4 for $\lambda$-almost all $u\in(0,1)$ that $(F_{Y}(Y)\mid X=q_{X}(u))$ and $(F_{Y^{\prime}}(Y^{\prime})\mid X^{\prime}=q_{X^{\prime}}(u))$ are not comparable in the dilation order (and thus neither in the dispersive order), see Figure 4. Hence, Proposition 5.2 cannot be applied to prove $\xi(Y,X)\geq\xi(Y^{\prime},X)\,.$ Nevertheless, the inequality for $\xi$ holds as a consequence of Theorem 1.8, where we have shown that $(Y,X)\succcurlyeq_{ccx}(Y^{\prime},X)$.
Figure 4: Left plot: Densities and means of the conditional distributions $F_{Y}(Y)|X=q_{X}(u)$ (solid) and $F_{Y^{\prime}}(Y^{\prime})|X^{\prime}=q_{X^{\prime}}(u)$ (dashed) in Example 5.5 for $\sigma=1$ and $\sigma^{\prime}=2$ and for u=0.1.u=0.1\,. The densities are determined by a kernel estimation based on a sample of size $10^{7}\,.$ Right plot: Densities from the left plot shifted by the mean. Since the supports are non-nested intervals satisfying (33), they are not comparable in convex order.
5.3 Variability of regression functions and magnitude of error curves
The global dependence orderings studied in [58, 59] compare conditional expectations and conditional variances. As we discuss in the sequel, they are not sufficient for ordering results on dependence measures like $\xi$.
Therefore, let $(Y,X)$ be a bivariate random vector. Then there exists a representation $Y=f(X,U)$ a.s. for a measurable function $f$ and a random variable $U$ that is uniform on $(0,1)$ and independent of $X$. The variable $U$ can be interpreted as noise. Informally, a weak (strong) influence of $U$ on $Y$ indicates strong (weak) dependence of $Y$ on $X$.
The idea in [58] is to define global dependence orderings for the regression function $m(x)=\mathbb{E}[Y|X=x]$ and for the error function $e(x)=\mathrm{Var}(Y|X=x)$. For ordering the magnitude of $e(X)$, the authors propose the (standard) stochastic order. For $m(X)$, they consider variability orderings such as the convex order, the dilation order, and the dispersive order. This approach comes closest to what we have been able to find in the literature on our approach, which compares the variability of the regression functions $P(Y\geq q_{Y}(v)\mid X)=\mathbb{E}[\mathds{1}_{Y\geq q_{Y}(v)}\mid X]$ for all $v\in(0,1)$. This is crucial for a dependence order underlying $\xi$.
In [59], the idea is to analyze the variability of the conditional expectation $\tilde{m}(u)=\mathbb{E}[Y|U=u]$ and the magnitude of the conditional variance $\tilde{e}(u)=\mathrm{Var}(Y|U=u)$. Similar to the approach in [58], this does not yield ordering results for dependence measures such as Chatterjee’s rank correlation.
Appendix
Appendix A Auxiliary results and preparation for proofs in Appendix B
For easier access to the proofs in Appendix B, we provide several useful results on generalized inverses, the convex order, the distributional transform, and the quantile transform.
A.1 Generalized inverses
For a distribution function $G$, we denote by $G^{-1}(v):=\inf\{y\mid G(y)\geq v\},$ $v\in(0,1)$, the left-continuous generalized inverse of $G$. We refer to [23] and [64, Chapter 21] for various properties of generalized inverses and will frequently use the following specific properties. Therefore, we denote by $G^{-}$ the left-continuous version of $G$ and we abbreviate $\iota_{G}:=G\circ G^{-1}$ and $\iota_{G}^{-}:=G^{-}\circ G^{-1}$.
Proposition A.1
  1. (i)
    $G^{-1}(v)\leq y$ if and only if $v\leq G(y)$;
  2. (ii)
    $\iota_{G}(t)=t=\iota_{G}^{-}(t)$ if and only if $G$ is continuous at $G^{-1}(t)$;
  3. (iii)
    $\iota_{G}^{-}$ is left-continuous while, in general, $\iota_{G}$ is neither left- nor right-continuous;
  4. (iv)
    $\iota_{G}(t)=t=\iota_{G}^{-}(t)$ for all $t\in(0,1)$ if and only if $G$ is continuous.
Denote by $y\mapsto\overline{F}_{Y}(y):=P(Y\geq y)$ the survival function of $Y$. Recall that we use the shorter notation $q_{Y}:=F_{Y}^{-1}$ for the generalized inverse of a random variable $Y$ with distribution function $F_{Y}$. The following example characterizes the marginal constraint in the definition of $\preccurlyeq_{ccx}$.
Lemma A.2 (Marginal constraint)
For real-valued random variables $Y$ and $Y^{\prime}$, the following statements are equivalent:
  1. (i)
    $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$;
  2. (ii)
    $\overline{F}_{Y}(q_{Y}(v))=\overline{F}_{Y^{\prime}}(q_{Y^{\prime}}(v))$ for all $v\in(0,1)$;
  3. (iii)
    $F_{Y}(q_{Y}(v))=F_{Y^{\prime}}(q_{Y^{\prime}}(v))$ for $\lambda$-almost all $v\in(0,1)$.
A.2 Convex order
For the proof of Theorem 1.4, we use of the following properties of convex order; see [47, Example 1.10.5] and [57, Section 3.A.2].
Lemma A.3 (Some properties of convex order)
Let $S_{n},T_{n},S,T$ be random variables with values in $[0,1]$. Then the following statements hold true:
  1. (i)
    Minimal elements: $\mathbb{E}S\leq_{cx}S$.
  2. (ii)
    Maximal elements: If $\mathbb{E}S=\mathbb{E}T$ and if $T$ takes values only in $\{0,1\}$, then $S\leq_{cx}T$.
  3. (iii)
    Antisymmetry: $S\leq_{cx}T$ and $S\geq_{cx}T$ implies $S\stackrel{{\scriptstyle\mathrm{d}}}{{=}}T$.
  4. (iv)
    Invariance under transformations: $S\leq_{cx}T$ implies $c\,S\leq_{cx}c\,T$ and $S+c\leq_{cx}T+c$ for all constants $c\in\mathbb{R}$.
  5. (v)
    Stability under weak convergence: If $S_{n}\xrightarrow[]{\penalty 10000\ d\penalty 10000\ }S$, $T_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }T$, $\mathbb{E}S_{n}\to\mathbb{E}S$, $\mathbb{E}T_{n}\to\mathbb{E}T$, and $S_{n}\leq_{cx}T_{n}$ for all $n$, then $S\leq_{cx}T$.
The conditional convex order is defined by comparing conditional survival probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order for all $v\in(0,1)$. Equivalent versions can be achieved for strict survival probabilities and cumulative probabilities as follows. Due to continuity properties of the transformations $\iota_{G}$ and $\iota_{G}^{-1}$ in Proposition A.1(ii) and (iii), exceptional null sets have to be considered.
Proposition A.4 (Versions of conditional convex order)
The following statements are equivalent:
  1. (i)
    $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$
  2. (ii)
    $P(Y<q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}<q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for all $v\in(0,1)$.
  3. (iii)
    $P(Y\leq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for $\lambda$-almost all $v\in(0,1)$.
  4. (iv)
    $P(Y>q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}>q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for $\lambda$-almost all $v\in(0,1)$.
The following example shows that the exceptional null set in the version of Proposition A.4(iii) is necessary.
Example A.5
Consider a random variable $Y$ with distribution function given by the left-hand graph in Figure 5. Let $Y^{\prime}=Y+2$ if $Y\leq 0.5\,$, and $Y^{\prime}=Y+1.5$ if Y>0.5.Y>0.5\,. Then, for $v=F_{Y}(0.5)=0.4\,$, we have $P(Y\leq q_{Y}(v))=0.4\neq 0.8=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$. Hence, $P(Y\leq q_{Y}(v)\mid\mathbb{X})$ and $P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ are not comparable in convex order, independent of the choice of $\mathbb{X}$. However, since $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ and since $F_{Y}^{-}\circ q_{Y}$ as well as $F_{Y^{\prime}}^{-}\circ q_{Y^{\prime}}$ are left-continuous, it follows that $P(Y\geq q_{Y}(v))=P(Y^{\prime}\geq q_{Y^{\prime}}(v))$ for all $v\in(0,1)$. Consequently, $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X})$ are comparable in conditionally convex order; more precisely, it is $(Y,\mathbb{X})=_{ccx}(F_{Y}(Y),\mathbb{X})=_{ccx}(F_{Y^{\prime}}(Y^{\prime}),\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X})$ for any choice of $\mathbb{X}$ using distributional invariance of $\preccurlyeq_{ccx}$ and $F_{Y^{\prime}}(Y^{\prime})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y}(Y)$.
Figure 5: Left: the distribution function of a random variable Y;Y\,; right: the distribution function of a random variable $Y^{\prime}$ which is a shift of $Y$ given by $Y^{\prime}=Y+2$ if $Y\leq 0.5\,$, and $Y^{\prime}=Y+1.5$ if Y>0.5;Y>0.5\,; see Example A.5. Since $Y$ and $Y^{\prime}$ have the same strength of functional dependence on a random vector $\mathbb{X}$ in terms of Chatterjee’s rank correlation (i.e., $\xi(Y|\mathbb{X})=\xi(Y^{\prime}|\mathbb{X})$), they should be equally ranked in a suitable global dependence order.
A.3 Distributional transform and quantile transform
Let $\mathbb{V}=(V_{1},\ldots,V_{p})$ be a random vector (on $(\Omega,\mathcal{A},P)$ which is assumed to be non-atomic) that is independent of $\mathbb{X}=(X_{1},\ldots,X_{p})$ and uniformly on $(0,1)^{p}$ distributed. Then the multivariate distributional transform $\tau_{\mathbb{X}}(\mathbb{X},\mathbb{V})$ of $\mathbb{X}$ (also known as generalized Rosenblatt transform) is defined by
$\tau_{\mathbb{X}}(\mathbb{x},\bm{\lambda}):=\big(F_{1}(x_{1},\lambda_{1}),F_{2}(x_{2},\lambda_{2}|x_{1})\dots,F_{p}(x_{p},\lambda_{p}|x_{1},\dots,x_{p-1})\big)$
for $\mathbb{x}=(x_{1},\ldots,x_{p})\in\mathbb{R}^{p}$ and all $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})\in[0,1]^{p}$, where
$F_{1}(x_{1},\lambda_{1})$ $:=$ $P(X_{1}<x_{1})+\lambda_{1}\,P(X_{1}=x_{1})$
$F_{i}(x_{i},\lambda_{i}|x_{1},\dots,x_{i-1})$ $:=$ $P(X_{i}<x_{i}|X_{1}=x_{1},\dots,X_{i-1}=x_{i-1})$
$+\,\lambda_{i}\,P(X_{i}=x_{i}|X_{1}=x_{1},\dots,X_{i-1}=x_{i-1})\,,\quad i\in\{2,\ldots,p\}\,;$
see [52, Section 1.3]. For $p=1$ and for a random variable $X$ with continuous distribution function $F_{X}\,$, the distributional transform $\tau_{X}(X,V)$ simplifies to $F_{X}(X)\,$, which is uniform on $(0,1)$.
As an inverse transformation of $\tau_{\mathbb{X}}\,$, we consider for a random vector $\mathbb{U}=(U_{1},\ldots,U_{p})$, uniformly on $(0,1)^{p}$ distributed, the multivariate quantile transform ${\bf q}_{\mathbb{X}}(\mathbb{U}):=(\zeta_{1},\ldots,\zeta_{p})$. It is iteratively defined by
$\begin{split}\zeta_{1}&:=F_{X_{1}}^{-1}(U_{1})\,,\\ \zeta_{i}&:=F^{-1}_{X_{i}|X_{i-1}=\xi_{i-1},\ldots,X_{1}=\xi_{1}}(U_{i})\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{for all }i\in\{2,\ldots,p\}\end{split}$ (36)
where $F_{W|\mathbb{Z}=\mathbb{z}}$ and $F^{-1}_{W|\mathbb{Z}=\mathbb{z}}$ denote the conditional distribution function of $W$ and its generalized inverse given $\mathbb{Z}=\mathbb{z}$.
According to [52, Theorem 1.12],
[…]
${\bf q}_{\mathbb{X}}(\mathbb{U})$ $\text{ is a random vector with distribution function }F_{\mathbb{X}},$ (38)
and the multivariate quantile transform is inverse to the multivariate distributional transform, i.e.,
$\mathbb{X}={\bf q}_{\mathbb{X}}\big(\tau_{\mathbb{X}}(\mathbb{X},\mathbb{V})\big)\qquad P\text{-almost surely.}$ (39)
Appendix B Proofs
B.1 Proof of Theorem 1.4
Axioms 1 and 2 follow from the definition of the conditional convex order (Definition 1.2).
To verify Axiom 3, first note that the deterministic random variable $P(Y\geq q_{Y}(v))$ is minimal in convex order within the class of $[0,1]$-valued random variables having mean $P(Y\geq q_{Y}(v))$; see Lemma A.3. Therefore, if $\mathbb{X}$ and $Y$ are independent, then $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ for all random vectors $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. For the reverse direction, assume that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ for all random vectors $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}\,.$ This implies for deterministic $\mathbb{X}^{\prime}$, say $\mathbb{X}^{\prime}=\mathbf{0}$ a.s., that
$P(Y\geq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\geq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})=P(Y^{\prime}\geq q_{Y^{\prime}}(v))=P(Y\geq q_{Y}(v))$
for all $v\in[0,1]$. Since the unconditional random variable $P(Y\geq q_{Y}(v))$ is deterministic and thus minimal in convex order (see Lemma A.3), we have for all $y\in\mathbb{R}$ that $P(Y\geq y\mid\mathbb{X})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}P(Y\geq y)$. This implies for all $y\in\mathbb{R}$ that $P(Y\geq y\mid\mathbb{X}=\mathbb{x})=P(Y\geq y)$ for $P^{\mathbb{X}}$-almost all $\mathbb{x}$, and thus $\mathbb{X}$ and $Y$ are independent.
To verify Axiom 4, first assume that $Y$ perfectly depends on $\mathbb{X}$, i.e., there exists a measurable function $f$ such that $Y=f(\mathbb{X})$ almost surely. This implies for all $v\in(0,1)$ that
$P(Y\geq q_{Y}(v)\mid\mathbb{X})$ $=\mathds{1}_{\{f(\mathbb{X})\geq q_{Y}(v)\}}\,.$
The random variable $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ is a maximal element in convex order within the class of $[0,1]$-valued random variables having mean $P(Y\geq q_{Y}(v))$; see Lemma A.3. Hence, for $(Y^{\prime}\mathbb{X}^{\prime})\in\mathcal{R}_{A}$, using the marginal constraint, it follows that $P(Y^{\prime}\geq q_{Y^{\prime}}(v))=P(Y\geq q_{Y}(v))$ and thus $P(Y^{\prime}\geq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})\leq_{cx}P(Y\geq q_{Y}(v)\mid\mathbb{X})$ for all $v\in(0,1)$. Therefore, $(Y^{\prime},\mathbb{X}^{\prime})\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all random vectors $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. For the reverse direction, assume that $(Y^{\prime},\mathbb{X}^{\prime})\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. This implies, in particular, $(Y,Y)\preccurlyeq_{ccx}(Y,\mathbb{X})$ and thus for all $v\in(0,1)$ that
$P(Y\geq q_{Y}(v)\mid\mathbb{X})\geq_{cx}P(Y\geq q_{Y}(v)\mid Y)=\mathds{1}_{\{Y\geq q_{Y}(v)\}}\,.$
The right-hand side in the above inequality is a random variable that is maximal in convex order within the class of $[0,1]$-valued random variables having mean $P(Y\geq q_{Y}(v))$ (see Lemma A.3). This implies equality in convex order and, by antisymmetry of convex order, for all $v\in(0,1)$ that $P(Y\geq q_{Y}(v)\mid\mathbb{X})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}P(Y\geq q_{Y}(v)\mid Y)=\mathds{1}_{\{Y\geq q_{Y}(v)\}}$. It follows that, for all $y\in\mathbb{R}$, $P(Y\geq y\mid\mathbb{X}=\mathbb{x})\in\{0,1\}$ for $P^{\mathbb{X}}$-almost all $\mathbb{x}$. Hence, the conditional distribution $Y\mid\mathbb{X}=\mathbb{x}$ is degenerate for $P^{\mathbb{X}}$-almost all $\mathbb{x}$, which means that $Y$ is a measurable function of $\mathbb{X}$; see the reasoning in [7, Proof of Theorem 9.2].
To verify the information gain inequality in Axiom 5, we have
$\mathbb{E}\varphi\left(P(Y\geq q_{Y}(v)\mid\mathbb{X})\right)$ $=\mathbb{E}\left[\mathbb{E}\varphi\left(P(Y\geq q_{Y}(v)\mid\mathbb{X})\right)\mid(\mathbb{X},\mathbb{Z})\right]$
$\leq\mathbb{E}\varphi\left(\mathbb{E}\left[P(Y\geq q_{Y}(v)\mid\mathbb{X})\mid(\mathbb{X},\mathbb{Z})\right]\right)$
$=\mathbb{E}\varphi\left(P(Y\geq q_{Y}(v)\mid(\mathbb{X},\mathbb{Z}))\right)$
using Jensen’s inequality for conditional expectation.
To verify Axiom 6, assume first that $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))\,.$ This implies $\xi(Y,\mathbb{X})=\xi(Y,(\mathbb{X},\mathbb{Z}))\,.$ The latter means that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}\,;$ see [7, Lemma 11.2]. For the reverse direction, conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$ implies for all $v$ that $P(Y\geq q_{Y}(v)\mid\mathbb{X})=P(Y\geq q_{Y}(v)\mid(\mathbb{X},\mathbb{Z}))$ almost surely and thus $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$.
To verify Axiom 7, observe that $\overline{\mathsf{Ran}(F_{g(Y)})}=\overline{\mathsf{Ran}(F_{Y})}$. It follows that $P(g(Y)\geq q_{g(Y)}(v))=P(Y\geq q_{Y}(v))$ for all $v\in(0,1)\,;$ see Lemma A.2. Hence, the invariance property for $Y$ follows from the definition of conditional probability and from the definition of conditional convex order. To show the invariance property for $\mathbb{X}$, let $\mathbb{h}$ be bijective. This implies that the $\sigma$-algebras generated by $\mathbb{X}$ and $\mathbb{h}(\mathbb{X})$ coincide. Hence, the statement follows from the definition of the conditional convex order.
Axiom 8: Since $\overline{\mathsf{Ran}(F_{F_{Y}(Y)})}=\overline{\mathsf{Ran}(F_{Y})}$, distributional invariance for $Y$ follows similar to the proof of Axiom 7. To prove distributional invariance for $\mathbb{X}$, it is straightforward to verify that $\xi(X_{k},F_{X_{k}}(X_{k}))=1$. Hence, there exists a measurable function $f_{k}$ such that $X_{k}=f_{k}(F_{X_{k}}(X_{k}))$ almost surely. Applying the data processing inequality twice, we obtain
$\big(Y,(F_{X_{1}}(X_{1}),\ldots,F_{X_{p}}(X_{p}))\big)$ $\preccurlyeq_{ccx}\big(Y,(X_{1},\ldots,X_{p})\big)$
$=_{ccx}\big(Y,(f_{1}(F_{X_{1}}(X_{1})),\ldots,f_{p}(F_{X_{p}}(X_{p})))\big)$
$\preccurlyeq_{ccx}\big(Y,(F_{X_{1}}(X_{1}),\ldots,F_{X_{p}}(X_{p}))\big),$
which proves the statement. $\blacksquare$
B.2 Proof of Theorem 1.6
The proof of Theorem 1.6 is based on Theorem 2.2 and Proposition 2.7 which we prove first.
The conditional convex order $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ is equivalent to $P(Y\leq q_{Y}(v)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ for $\lambda$-almost all $v\in(0,1)$; see Proposition A.4. However, the latter is by the Hardy-Littlewood-Polya theorem (see Lemma 2.1) equivalent to $\eta_{Y|\mathbb{X}}^{v}\prec_{S}\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v}$ for $\lambda$-almost all $v\in(0,1)$. $\blacksquare$
For the proof of Proposition 2.7, we make use of the following result on rearrangements of conditional distribution functions.
Lemma B.1 (Rearranged conditional distribution function)
Let $(Y,\mathbb{X})$ be a $(1+p)$-dimensional random vector. Then there exists a function $\mathbb{R}\times(0,1)\ni(y,u)\mapsto F_{u}(y)$ such that $u\mapsto F_{u}(y)$ is a decreasing rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$ for all $y\in\mathbb{R}$ and $y\mapsto F_{u}(y)$ is a distribution function for all $u\in(0,1)$.
For $y\in\mathbb{R}$, denote by $u\mapsto G_{u}(y)$ a decreasing rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. Since $P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))\to 0$ for all $\mathbb{u}$ as $y\to$-$\infty$, also $G_{u}(y)\to 0$ for all $u$ as $y\to$-$\infty$. Similarly, $G_{u}(y)\to 1$ for all $u$ as $y\to+\infty$. For $y\leq y^{\prime}$, it is $P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))\leq P(Y\leq y^{\prime}\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$ for all $\mathbb{u}$. Hence, also the decreasing rearrangements satisfy $G_{u}(y)\leq G_{u}(y^{\prime})$ for all $u$. Define $H(u,y):=\int_{0}^{u}G_{t}(y)\mathsf{\,d}t$ for $(u,y)\in(0,1)\times\mathbb{R}$. It is straightforward to verify that $H$ is a bivariate distribution function. Then, by the regular version of conditional probability [33, Theorem 6.3], there exists a family $(F_{u})_{u\in[0,1]}$ of distribution functions such that
$\int_{0}^{u}F_{t}(y)\mathsf{\,d}t=H(u,y)=\int_{0}^{u}G_{t}(y)\mathsf{\,d}t\quad\text{for all }(u,y)\in(0,1)\times\mathbb{R}.$
For fixed $y$, it follows that $F_{t}(y)=G_{t}(y)$ for $\lambda$-almost all $t\in(0,1)$. This implies
$\lambda(\{t\mid F_{t}(y)\geq w\})=\lambda(\{t\mid G_{t}(y)\geq w\})\quad\text{for all }w\in\mathbb{R}.$
Hence, also $F_{u}$ is a decreasing rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. Since $F_{u}$ is a distribution function for all $u$, the statement is proven. $\blacksquare$
To show statement (i), we obtain for all $y\in\mathbb{R}$ that
$P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq y\big)$ $=\int_{[0,1]}P(F_{u}^{-1}(V)\leq y\mid U=u)\mathsf{\,d}\lambda(u)=\int_{[0,1]}P(F_{u}^{-1}(V)\leq y)\mathsf{\,d}\lambda(u)$
$=\int_{[0,1]}F_{u}(y)\mathsf{\,d}\lambda(u)=\int_{[0,1]^{p}}P\big(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u})\big)\mathsf{\,d}\lambda^{p}(\mathbb{u})$
$=\int_{\Omega}P(Y\leq y\mid\mathbb{X})\mathsf{\,d}P=P(Y\leq y),$
where the use for the first equality the definition of the rearranged quantile transform in (10). For the second equality, we use independence of $U$ and $V$. The third equality follows from Proposition A.1(i), using the fact that $F_{u}$ is a distribution function, and from the fact that $V$ is uniform on $(0,1)$. The fourth equality holds by definition of $F_{u}$ as a rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. The fifth equality follows with (38), and the last equality is due to disintegration.
To show statement (ii), recall that the family $\{q_{Y|\mathbb{X}}^{\uparrow u}(V)\}_{[0,1]}$ is SI in $u$, that is, $\mathbb{E}f(q_{Y|\mathbb{X}}^{\uparrow u}(V))=\mathbb{E}[f(q_{Y|\mathbb{X}}^{\uparrow u}(V))\mid U=u]$ is increasing in $u$ for all increasing functions $f$ such that the expectations exist, where we use for the equality that $U$ and $V$ are independent. However, the latter means that $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ is a bivariate SI random vector.
To show statement (iii), let $v\in(0,1)$. Then we obtain for all $u\in[0,1]$ that
$\int_{0}^{u}P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{q_{Y|\mathbb{X}}^{\uparrow U}}(v)\mid U=s\big)\mathsf{\,d}s$ $=\int_{0}^{u}P\big(q_{Y|\mathbb{X}}^{\uparrow s}(V)\leq q_{Y}(v)\big)\mathsf{\,d}s$
$=\int_{0}^{u}F_{s}(q_{Y}(v))\mathsf{\,d}s=\int_{0}^{u}\big(\eta_{Y|\mathbb{X}}^{v}\big)^{*}(s)\mathsf{\,d}s,$
where we use for the first equality $q_{Y|\mathbb{X}}^{\uparrow U}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and independence of $U$ and $V$. The second equality follows from (10) and Proposition A.1(i). The last equality is due to (9). Since the integrand of the first integral is decreasing in $s$, it coincides $\lambda$-almost surely with the decreasing rearrangement $(\eta_{Y|\mathbb{X}}^{v})^{*}$. Hence, the characterization of $\preccurlyeq_{ccx}$ through the Schur order in Theorem 2.2 implies $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})$. $\blacksquare$
We also use the following result, which gives several variants of the concordance order. Recall that $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$ is defined by $P(Y\leq y,X\leq x)\leq P(Y^{\prime}\leq y,X^{\prime}\leq x)$ and $P(Y>y,X>x)\leq P(Y^{\prime}>y,X^{\prime}>x)$ for all $x,y\in\mathbb{R}$.
Proposition B.2 (Concordance order)
Let $X,Y,X^{\prime},Y^{\prime}$ be real-valued random variables. Then the following statements are equivalent:
  1. (i)
    $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$,
  2. (ii)
    $P(Y\leq y,X\leq x)\leq P(Y^{\prime}\leq y,X^{\prime}\leq x)$ and $P(Y\geq y,X\geq x)\leq P(Y^{\prime}\geq y,X^{\prime}\geq x)$ for all $x,y\in\mathbb{R}$.
  3. (iii)
    $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$, $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$, and $P(Y\leq y,X\leq x)\leq P(Y^{\prime}\leq y,X^{\prime}\leq x)$ for all $x,y\in\mathbb{R}$.
  4. (iv)
    $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$, $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$, and $P(Y\geq y,X\geq x)\leq P(Y^{\prime}\geq y,X^{\prime}\geq x)$ for all $x,y\in\mathbb{R}$.
’(i) $\Longrightarrow$ (ii)’: For $x_{n}\uparrow x$ and $y_{n}\uparrow y$, continuity of measures yields
$P(Y\geq y,X\geq x)$ $=\lim_{n\to\infty}P(Y>y_{n},X>x_{n})$
$\leq\lim_{n\to\infty}P(Y^{\prime}>y_{n},X^{\prime}>x_{n})=P(Y^{\prime}\geq y,X^{\prime}\geq x).$
’(ii) $\Longrightarrow$ (i)’: For $x_{n}\downarrow x$ and $y_{n}\downarrow y$, continuity of measures yields
$P(Y>y,X>x)$ $=\lim_{n\to\infty}P(Y\geq y_{n},X\geq x_{n})$
$\leq\lim_{n\to\infty}P(Y^{\prime}\geq y_{n},X^{\prime}\geq x_{n})=P(Y^{\prime}>y,X^{\prime}>x).$
’(i) $\Longrightarrow$ (iii)’: The definition of the concordance order implies $P(X\leq x)\leq P(X^{\prime}\leq x)$ and $P(X>x)\leq P(X^{\prime}>x)$ for all $x$. This yields $P(X\leq x)\leq P(X^{\prime}\leq x)$ and $P(X\leq x)\geq P(X^{\prime}\leq x)$ for all $x$ and thus $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$. Similarly, it follows that $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$.
’(iii) $\Longrightarrow$ (i)’: By the inclusion-exclusion principle, we have
$P(X>x,Y>y)$ $=1-P(X\leq x)-P(Y\leq y)+P(X\leq x,Y\leq y)$
$\leq 1-P(X^{\prime}\leq x)-P(Y^{\prime}\leq y)+P(X^{\prime}\leq x,Y^{\prime}\leq y)$
$=P(X^{\prime}>x,Y^{\prime}>y).$
’(i) $\Longrightarrow$ (iv)’: Similar as above, we obtain $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$. Further, for $x_{n}\uparrow x$ and $y_{n}\uparrow y$, it follows that
$P(Y\geq y,X\geq x)$ $=\lim_{n\to\infty}P(Y>y_{n},X>x_{n})$
$\leq\lim_{n\to\infty}P(Y^{\prime}>y_{n},X^{\prime}>x_{n})=P(Y^{\prime}\geq y,X^{\prime}\geq x).$
’(iv) $\Longrightarrow$ (i)’: Follows similar as above using an approximation of measures and the inclusion-exclusion principle. $\blacksquare$
We also need the following variants of the pointwise order of distribution functions under the marginal constraint.
Lemma B.3
Let $Y,Y^{\prime}$ be real-valued random variables, and let $U$ be uniform on $(0,1).$ Under the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$, the following statements are equivalent:
  1. (i)
    $P(Y\leq q_{Y}(v),U\leq u)\leq P(Y^{\prime}\leq q_{Y^{\prime}}(v),U\leq u)$ for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1),$
  2. (ii)
    $P(F_{Y}(Y)\geq v,U\geq u)\leq P(F_{Y^{\prime}}(Y^{\prime})\geq v,U\geq u)$ for all $(u,v)\in(0,1)^{2}$.
Assume (i). Denote by $N_{1}\subset[0,1]$ the exceptional $\lambda$-null set such that $P(Y\leq q_{Y}(v),U\leq u)\leq P(Y^{\prime}\leq q_{Y^{\prime}}(v),U\leq u)$ for all $v\in(0,1)\setminus N_{1}$ and $u\in(0,1)$. Noting that the marginal constraint implies $P(Y\leq q_{Y}(v))=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$ for all $v\in[0,1]$ outside a $\lambda$-null set $N_{2}$ (see Lemma A.2), we obtain from the inclusion-exclusion principle that
$P(Y>q_{Y}(v),U>u)\leq P(Y^{\prime}>q_{Y^{\prime}}(v),U>u)$ (40)
for all $v\in(0,1)\setminus N$ and for all $u\in(0,1)$, where $N:=N_{1}\cup N_{2}$. We aim show to that
$P(Y\geq q_{Y}(v),U\geq u)\leq P(Y^{\prime}\geq q_{Y^{\prime}}(v),U\geq u)$ (41)
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Then, using the identities $\{Y\geq q_{Y}(v)\}=\{F_{Y}(Y)\geq v\}$ and $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}=\{F_{Y^{\prime}}(Y^{\prime})\geq v\}$ (see Proposition A.1(i)), we obtain
$P(F_{Y}(Y)\geq v,U\geq u)\leq P(F_{Y^{\prime}}(Y^{\prime})\geq v,U\geq u)$ (42)
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Finally, continuity of measures from above yields statement (ii). To prove (41), we consider two cases. For $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})\setminus N$, choose $v_{n}\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})\setminus N$ with $v_{n}\uparrow v$. Then we have $q_{Y}(v_{n})\uparrow q_{Y}(v)$ and, using the marginal constraint, we also obtain $q_{Y^{\prime}}(v_{n})\uparrow q_{Y^{\prime}}(v)$. Hence, by continuity of measures from above, we obtain with $u_{n}\uparrow u$ that
$\begin{split}P(Y\geq q_{Y}(v),U\geq u)&=\lim_{n\to\infty}P(Y>q_{Y}(v_{n}),U>u_{n})\\ &\leq\lim_{n\to\infty}P(Y^{\prime}>q_{Y^{\prime}}(v_{n}),U>u_{n})=P(Y^{\prime}\geq q_{Y^{\prime}}(v),U^{\prime}\geq u),\end{split}$ (43)
where the inequality follows from (40). For the second case, let $v\in(0,1)\setminus(\overline{\mathsf{Ran}(F_{Y})})\cup N)$ and define $v_{0}:=F_{Y}^{-}(q_{Y}(v))$. Note that, by the marginal constraint and by left-continuity of the transformation $\iota_{F_{Y}}^{-}=F_{Y}^{-}\circ F_{Y}^{-1}$ in Proposition A.1(iii), we also have $v_{0}=F_{Y^{\prime}}^{-}(q_{Y^{\prime}}(v))$. Now, choose $v_{n}\in[0,1]\setminus N$ with $v_{n}\uparrow v_{0}$. Then
$\{Y\geq q_{Y}(v)\}=\{Y\geq q_{Y}(v_{0})\}=\bigcap_{n\in\mathbb{N}}\{Y>q_{Y}(v_{n})\}\cap\{Y\notin[q_{Y}(v_{0}),q_{Y}(v))\}.$ (44)
Using that $\{Y\in[q_{Y}(v_{0}),q_{Y}(v))\}$ is a $P$-null set and using a similar reasoning for $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}$, we obtain (43) also in the second case. This proves (41) and thus statement (ii).
For the reverse direction, assume that (42) holds for all $(u,v)\in(0,1)^{2}$. Using $\{Y\geq q_{Y}(v)\}=\{F_{Y}(Y)\geq v\}$ and $\{Y^{\prime}\geq q_{Y^{\prime}}(v)\}=\{F_{Y^{\prime}}(Y^{\prime})\geq v\}$, we obtain (41) for all $(u,v)\in(0,1)^{2}$. We aim to show (40) for $\lambda$-almost all $v\in(0,1)$ and for all $u\in[0,1]$, which implies statement (i). To this end, consider first the case where $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})$. Choose $v_{n}\downarrow v$ and $u_{n}\downarrow u$. Then we have $\bigcap_{n\in\mathbb{N}}\{Y\geq q_{Y}(v_{n})\}=\{Y>q_{Y}(v)\}\setminus\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ for $q_{Y}^{+}(v)=\sup\{y\mid F_{Y}(y)\leq v\}$. Since $\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ is a $P$-null set and due to a similar reasoning for $Y^{\prime}$ (using the marginal constraint), we obtain (40). For the second case, let $v\in(0,1)\setminus\overline{\mathsf{Ran}(F_{Y})}$ and define $v_{0}:=F_{Y}(q_{Y}(v))=F_{Y^{\prime}}(q_{Y^{\prime}}(v))$. Then we have $\bigcap_{n\in\mathbb{N}}\{Y\geq q_{Y}(v_{n})\}=\{Y>q_{Y}(v_{0})\}\setminus N_{v_{0}}$ for some $Y$-null set $N_{v_{0}}$. Using a similar reasoning for an approximation of $\{Y^{\prime}>q_{Y^{\prime}}(v)\}$ from above, we obtain (40) from (41). $\blacksquare$
We are now able to prove Theorem 1.6.
First, recall that a comparison in the concordance order implies identical marginal distributions; see Proposition B.2. Thus (ii) implies
$F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\,.$ (45)
Second, $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ which is equivalent to
$F_{Y}\circ q_{Y}(v)=F_{Y^{\prime}}\circ q_{Y^{\prime}}(v)\quad\text{for }\lambda\text{-almost all }v\in(0,1);$ (46)
see Lemma A.2. Since $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $Y^{\prime}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)$ due to Proposition 2.7(i), we obtain that (45) and (46) are equivalent.
Now, to show ’(i) $\Longrightarrow$ (ii)’, assume that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$ Then we obtain
$\begin{split}P(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u)&=\int_{0}^{u}P\left(q^{\uparrow U}_{Y|\mathbb{X}}(V)\leq q_{Y}(v)\mid U=s\right)\mathsf{\,d}s\\ &=\int_{0}^{u}P\left(q^{\uparrow s}_{Y|\mathbb{X}}(V)\leq q_{Y}(v)\right)\mathsf{\,d}s\\ &=\int_{0}^{u}P\left(V\leq F_{s}(q_{Y}(v))\right)\mathsf{\,d}s\\ &=\int_{0}^{u}(\eta_{Y|\mathbb{X}}^{v})^{*}(s)\mathsf{\,d}s\\ &\leq\int_{0}^{u}(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v})^{*}(s)\mathsf{\,d}s\\ &=\cdots=P(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u)\end{split}$ (47)
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in[0,1]$ with equality for $u=1,$ where we use for the inequality the characterization of $\preccurlyeq_{ccx}$ by the Schur order; see Theorem 2.2. The first equality is due to the disintegration theorem, the second equality holds by independence of $U$ and $V.$ For the third equality, we apply Proposition A.1(i) using that $F_{s}$ is a distribution function with associated quantile function $q_{Y|\mathbb{X}}^{\uparrow s}$; see (10). The fourth equality follows with (9) noting that $V$ is uniform on $(0,1)$.
Using again $q_{Y|\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and $q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and using the marginal constraint, we apply Lemma B.3 to prove that (47) implies
$P\big(F_{Y}(q_{Y|\mathbb{X}}^{\uparrow U}(V))\geq v,U\geq u\big)\leq P\big(F_{Y^{\prime}}(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V))\geq v,U\geq u\big)$ (48)
for all $(u,v)\in(0,1)^{2}$. Now, statement (ii) follows from (48) by Proposition B.2 where we use the identity in (45).
To prove the reverse direction, assume that $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$. Applying Proposition B.2, we obtain (45) and the inequality (48) for all $(u,v)\in[0,1]^{2}$. Using that (45) is equivalent to (46) which implies the marginal constraint, we obtain from (48) by Lemma B.3 that
$P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u\big)\leq P\big(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u\big)$ (49)
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$, where $U$ and $V$ are independent and uniform on $(0,1)$. By a similar reasoning as in (47), it follows that
$\int_{0}^{u}(\eta_{Y|\mathbb{X}}^{v})^{*}(s)\mathsf{\,d}s$ $=P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{Y}(v),U\leq u\big)$
$\leq P\big(q_{Y^{\prime}|\mathbb{X}^{\prime}}^{\uparrow U}(V)\leq q_{Y^{\prime}}(v),U\leq u\big)=\int_{0}^{u}(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v})^{*}(s)\mathsf{\,d}s$
for $\lambda$-almost all $v\in(0,1)$ and for all $u\in(0,1)$. Due to the characterization of the conditional convex order by the Schur order in Theorem 2.2, we obtain $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. $\blacksquare$
B.3 Proof of Theorem 1.8
The proof of Theorem 1.8 is based on Theorem 2.2, which we have already proven, and on Theorem 3.3, which we prove now.
(i): Due to sufficiency of $g$ for $Y|\mathbb{X}$ and of $h$ for $Y^{\prime}|\mathbb{X}^{\prime}\,$, we obtain from self-equitability of $\preccurlyeq_{ccx}$ (see Remark 1.5(d)) that $(Y,\mathbb{X})=_{ccx}(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$
(ii): Since $Y\uparrow_{st}g(\mathbb{X})$ and $Y^{\prime}\uparrow_{st}h(\mathbb{X}^{\prime})\,$, the (on $\mathsf{Ran}(F_{Y})\times\mathsf{Ran}(F_{g(\mathbb{X})})$ resp. $\mathsf{Ran}(F_{Y^{\prime}})\times\mathsf{Ran}(F_{h(\mathbb{X})})$ uniquely determined) copulas $C_{Y,g(\mathbb{X})}$ and $C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ can chosen to be SI on $0,1]^{2}$. Hence, by Theorem 1.6, $C_{Y,g(\mathbb{X})}\leq_{c}C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ and $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}}})$ imply $(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))\,.$ Then, the statement follows from part (i).
Statement (iii) follows from (ii) replacing $g(\mathbb{X})$ and $h(\mathbb{X}^{\prime})$ by $g(\mathbb{X})$ and $h(\mathbb{X}^{\prime})\,. \blacksquare$
For the proof of Theorem 1.8, we also need the following simple lemma, whose proof is straightforward.
Lemma B.4 (Sign change criterion)
Let $f,g\colon(0,1)\to\mathbb{R}$ be decreasing, Lebesgue-integrable functions with $\int f\mathsf{\,d}\lambda=\int g\mathsf{\,d}\lambda\,.$ Assume that there exists $t_{*}\in(0,1)$ such that $\{f>g\}\subseteq[0,t_{*}]\subseteq\{f\geq g\}\,.$ Then we have $\int_{0}^{x}g(t)\mathsf{\,d}t\leq\int_{0}^{x}f(t)\mathsf{\,d}t$ for all $x\in(0,1)\,.$
For $\sigma=0\,$, the statement is trivial. Hence, assume $\sigma>0\,.$ For $Z:=f(\mathbb{X})$ and for $y,y^{\prime}\in\mathbb{R}\,$, we have
$P(Y\leq y\mid Z=q_{Z}(t))=P(Z+\sigma\varepsilon\leq y\mid Z=q_{Z}(t))=F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})\,$, (50)
where we use for the second equality independence of $\varepsilon$ and $f(\mathbb{X})\,.$ Similarly, it holds that $P(Y^{\prime}\leq y^{\prime}\mid Z=q_{Z}(t))=F_{\varepsilon}(\tfrac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}})\,.$ It follows with Proposition A.1 (i) that
$t$ $\leq F_{Z}(\tfrac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}})$ (51)
$\Longleftrightarrow$ $q_{Z}(t)$ $\leq\frac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}}$ (52)
$\Longleftrightarrow$ $\frac{y-q_{Z}(t)}{\sigma}$ $\geq\frac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}$ (53)
$\Longrightarrow$ $F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})$ $\geq F_{\varepsilon}(\tfrac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}),$ (54)
where we use $\sigma<\sigma^{\prime}$ for the second equivalence. Similarly, we have that $t>F_{Z}(\tfrac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}})$ implies $F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})\leq F_{\varepsilon}(\tfrac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}).$ Using the marginal constraint $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})},$ we obtain $P(Y\leq q_{Y}(v))=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$ for $\lambda$-almost all $v\in(0,1)\,.$ Hence, since $F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})$ and $F_{\varepsilon}(\tfrac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}})$ are decreasing in t,t\,, it follows with Lemma B.4 for $\lambda$-almost all $v\in(0,1)$ with $y:=q_{Y}(v)$ and $y^{\prime}:=q_{Y^{\prime}}(v)$ that
$\int_{0}^{x}P(Y\leq q_{Y}(v)\mid Z=q_{Z}(t))\mathsf{\,d}t\geq\int_{0}^{x}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid Z=q_{Z}(t))\mathsf{\,d}t\quad\text{for all }x\in[0,1]$ (55)
with equality for x=1.x=1\,. Using Theorem 2.2, it follows that $(Y,f(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},f(\mathbb{X}))\,.$ Since $Y\mid f(\mathbb{X})$ is sufficient for $Y\mid\mathbb{X}$ and, similarly, $Y^{\prime}\mid f(\mathbb{X})$ is sufficient for $Y^{\prime}\mid\mathbb{X}$, we obtain from Theorem 3.3(i) that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. $\blacksquare$
B.4 Remaining proofs of Section 2
The equivalence of (i) and (ii) is explained in detail in Example 2.4 and Figure 1. The condition $q=q^{\prime}$ thereby equals the marginal constraint.
(iii): $X$ and $Y$ are independent if and only if $p_{00}=(1-p)(1-q)$ if and only if $p_{01}=p(1-q)$ if and only if $\alpha=\beta$.
(iv): $Y$ perfectly depends on $X$ if and only if (1) $p_{01}=0$, $p=q$, and $p_{00}=1-p$ or (2) $p_{00}=0$, p=1qp=1-q, and $p_{01}=p$ if and only if $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$.
(v): $X$ and $Y$ are comonotone if and only if $p_{00}=(1-p)\wedge(1-q)$ if and only if $\alpha=\frac{1-q}{1-p}\wedge 1$.
(vi): $X$ and $Y$ are countermonotone if and only if $p_{00}=((1-p)+(1-q)-1)\vee 0$ if and only if $\alpha=0\vee\left(1-\frac{q}{1-p}\right)$. $\blacksquare$
B.5 Remaining proofs of Section 3
For $S$ defined in Example 2.9, a reduced random vector of $(Y,\mathbb{X})$ is given by $(F_{Y}(Y),\Phi(S))$; see (20). Hence, for $S^{\prime}$ defined analogously, we obtain from Theorem 1.6 that
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\Longleftrightarrow\quad(F_{Y}(Y),\Phi(S))\leq_{c}(F_{Y^{\prime}}(Y^{\prime}),\Phi(S^{\prime})).$ (56)
Since $(Y,S)$ is bivariate normal with $\mathrm{Var}(S)=1$, the distribution function of the random vector $(F_{Y}(Y),\Phi(S))$ is the Gaussian copula with correlation parameter $\varrho=\sqrt{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}}\,;$ see (14). Similarly, $(F_{Y^{\prime}}(Y^{\prime}),\Phi(S^{\prime}))$ follows a Gaussian copula with parameter $\varrho^{\prime}=\sqrt{\Sigma_{Y^{\prime},\mathbb{X}^{\prime}}\Sigma_{\mathbb{X}^{\prime}}^{-}\Sigma_{\mathbb{X}^{\prime},Y^{\prime}}/\sigma_{Y^{\prime}}^{2}}$. Since the Gaussian copula is increasing in its parameter with respect to the concordance order [3], the right-hand side of (56) is equivalent to $\varrho\leq\varrho^{\prime}\,.$ $\blacksquare$
B.6 Proofs of Section 4
We first prove that $\xi_{\varphi}$ is increasing in the conditional convex order. Due to the marginal constraint, $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies that the normalizing constants of $\xi_{\varphi}(Y,\mathbb{X})$ and $\xi_{\varphi}(Y^{\prime},\mathbb{X}^{\prime})$ are equal. Then, by the Hardy-Littlewood-Polya theorem (Lemma 2.1) and the characterization of $\preccurlyeq_{ccx}$ by the Schur order in Theorem 2.2, it follows that
$\alpha_{\varphi}\,\xi_{\varphi}(Y,\mathbb{X})$ $=\int_{\mathbb{R}}\int_{\mathbb{R}^{p}}\varphi(F_{Y|\mathbb{X}=\mathbb{x}}(y)-F_{Y}(y))\mathsf{\,d}P^{\mathbb{X}}(\mathbb{u})\mathsf{\,d}P^{Y}(y)$
$=\int_{(0,1)}\int_{(0,1)^{p}}\varphi\big(F_{Y|\mathbb{X}={\bf q}_{\mathbb{X}}(\mathbb{u})}(q_{Y}(v))-F_{Y}(q_{Y}(v))\big)\mathsf{\,d}\lambda^{p}(\mathbb{u})\mathsf{\,d}\lambda(v)$
$=\int_{(0,1)}\int_{(0,1)^{p}}\varphi(\eta_{Y|\mathbb{X}}^{v}(\mathbb{u})-F_{Y}(q_{Y}(v)))\mathsf{\,d}\lambda^{p}(\mathbb{u})\mathsf{\,d}\lambda(v)$
$\leq\int_{(0,1)}\int_{(0,1)^{p^{\prime}}}\varphi(\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v}(\mathbb{u})-F_{Y^{\prime}}(q_{Y^{\prime}}(v)))\mathsf{\,d}\lambda^{p^{\prime}}(\mathbb{u})\mathsf{\,d}\lambda(v)$
$=\cdots=\alpha_{\varphi}\,\xi_{\varphi}(Y^{\prime},\mathbb{X}^{\prime})$
where we use that $z\mapsto\varphi(z-c)$ is convex whenever $\varphi$ is convex and that $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. By a similar reasoning, $\Lambda_{\varphi}$ is increasing in $\preccurlyeq_{ccx}$.
To prove assertion (i), we use that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies (25) as already shown. Since the conditional convex order satisfies Axiom 3, $(Y,\mathbb{X})$ is a minimal element in $\preccurlyeq_{ccx}$ if $Y$ and $\mathbb{X}$ are independent. In this case, we have $\xi_{\varphi}(Y,\mathbb{X})=0=\Lambda_{\varphi}(Y,\mathbb{X})$. Similarly, the conditional convex order satisfies Axiom 4, and thus $(Y,\mathbb{X})$ is a maximal element in $\preccurlyeq_{ccx}$ if $Y$ perfectly depends on $\mathbb{X}$. In this case, we have $\xi_{\varphi}(Y,\mathbb{X})=1=\Lambda_{\varphi}(Y,\mathbb{X})$, which is also immediate from the definition of the functionals.
The zero-independence property in (ii) follows from the fact that $Y$ and $\mathbb{X}$ are independent if and only if $F_{Y|\mathbb{X}=\mathbb{x}}=F_{Y}$ for $P^{\mathbb{X}}$-almost every $\mathbb{x}\in\mathbb{R}^{p}$. This condition is equivalent to requiring that the inner integrals in the definitions of $\xi_{\varphi}$ and $\Lambda_{\varphi}$ vanish $P^{\mathbb{X}}$-almost everywhere and $P^{\mathbb{X}}\otimes P^{\mathbb{X}}$-almost everywhere, respectively, since $\varphi$ is strictly convex at 0 with $\varphi(0)=0$.
The max-functionality property of $\Lambda_{\varphi}$ is a direct extension of [5, Proof of Theorem 3.2] to a multivariate vector $\mathbb{X}=(X_{1},\ldots,X_{p})$. To prove max-functionality for $\xi_{\varphi}$, it remains to show that $\xi_{\varphi}(Y,\mathbb{X})=1$ implies $Y=f(\mathbb{X})$ almost surely. Therefore, define $f_{y}(\mathbb{x}):=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ and $a:=F_{Y}(y)$ and set
$c_{1-a}$ $:=f_{y}(\mathbb{x})-\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}\in[0,1]$
$c_{-a}$ $:=(1-f_{y}(\mathbb{x}))-\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}\in[0,1]$
$c_{b}$ $:=2\,\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}\in[0,1]\,.$
Then $c_{1-a}+c_{-a}+c_{b}=1$ and convexity of $\varphi$ gives
$\varphi(f_{y}(\mathbb{x})-a)$ $=\varphi\left((1-a)\,c_{1-a}+(-a)\,c_{-a}+\frac{1-2a}{2}c_{b}\right)$
$\leq\varphi(1-a)\,c_{1-a}+\varphi(-a)\,c_{-a}+\varphi\left(\frac{1-2a}{2}\right)\,c_{b}$
$=\varphi(1-a)\,f_{y}(\mathbb{x})+\varphi(-a)\,(1-f_{y}(\mathbb{x}))$
$\qquad-\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}\,\left(\underbrace{\varphi(1-a)+\varphi(-a)-2\,\varphi\left(\frac{1-2a}{2}\right)}_{=:c(a)}\right)$
$=\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-a\right)\mathsf{\,d}P^{Y|\mathbb{X}=\mathbb{x}}(y_{1})-\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}\,c(a)\,$,
where $c(a)>0$ follows from strict convexity of $\varphi$ at 0 with $\varphi(0)=0$. Consequently,
$1$ $=\xi_{\varphi}(Y,\mathbb{X})=\alpha_{\varphi}^{-1}\,\int_{\mathbb{R}}\int_{\mathbb{R}^{p}}\varphi(f_{y}(\mathbb{x})-a)\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})\mathsf{\,d}P^{Y}(y)$
$\leq\alpha_{\varphi}^{-1}\,\int_{\mathbb{R}}\int_{\mathbb{R}^{p}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-a\right)\mathsf{\,d}P^{Y|\mathbb{X}=\mathbb{x}}(y_{1})\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})\mathsf{\,d}P^{Y}(y)$
$\qquad-\alpha_{\varphi}^{-1}\,\int_{\mathbb{R}}\int_{\mathbb{R}^{p}}\underbrace{c(a)}_{>0}\cdot\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x})\mathsf{\,d}P^{Y}(y)$
$\leq\alpha_{\varphi}^{-1}\,\int_{\mathbb{R}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-a\right)\mathsf{\,d}P^{Y}(y_{1})\mathsf{\,d}P^{Y}(y)=1\,.$
Hence, there exists some Borel set $B\subseteq\mathbb{R}^{p}$ with $P^{\mathbb{X}}(B)=1$ such that $\min\{f_{y}(\mathbb{x}),1-f_{y}(\mathbb{x})\}=0$ for all $\mathbb{x}\in B$, implying that $f_{y}(\mathbb{x})=F_{Y|\mathbb{X}=\mathbb{x}}(y)\in\{0,1\}$ for all $\mathbb{x}\in B$. Now, proceeding as in the proof of Theorem 3.2 in [5] yields the claimed max-functionality property of $\xi_{\varphi}$.
For verifying (iii), it remains to show that $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ implies conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$. Therefore, define again $f_{y}(\mathbb{x}):=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ and $g_{y}(\mathbb{x},\mathbb{z}):=F_{Y|\mathbb{X}=\mathbb{x},\mathbb{Z}=\mathbb{z}}(y)\,.$ Using $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ and Jensen’s inequality, it follows that
$\beta_{\varphi}\Lambda_{\varphi}(Y,\mathbb{X})=\int\int\varphi(f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2}))\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}_{1})\otimes P^{\mathbb{X}}(\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$=\int\int\varphi\left(\int g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})-g_{y}(\mathbb{x}_{2},\mathbb{z}_{2})\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{1}}(\mathbb{z}_{1})\otimes P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{2}}(\mathbb{z}_{2})\right)\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}_{1})\otimes P^{\mathbb{X}}(\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$\leq\int\int\int\varphi(g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})-g_{y}(\mathbb{x}_{2},\mathbb{z}_{2}))\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{1}}(\mathbb{z}_{1})\otimes P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{2}}(\mathbb{z}_{2})\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}_{1})\otimes P^{\mathbb{X}}(\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$=\beta_{\varphi}\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))=\beta_{\varphi}\Lambda_{\varphi}(Y,\mathbb{X})\,$,
and thus the inequality becomes an equality. Define $\mathbb{Z}_{\mathbb{x}}^{i}:=q_{\mathbb{Z}|\mathbb{X}=\mathbb{x}}(\mathbb{U}_{i})$ for $i\in\{1,2\}$ and $\mathbb{U}_{1},\mathbb{U}_{2}$ i.i.d. uniformly distributed on $0,1]^{r},$ where $r$ is the dimension of $\mathbb{Z}$. Hence, there exist a $P^{Y}$-null set $N$ and for each $y\in N^{c}$ some Borel set $G_{y}$ with $P^{\mathbb{X}}\otimes P^{\mathbb{X}}(G_{y})=1$ such that
$\varphi(f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2}))=\mathbb{E}\varphi\left(g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})\right)$ (57)
for all $(\mathbb{x}_{1},\mathbb{x}_{2})\in G_{y}$ and $y\in N^{c}\,.$ It follows that
$g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})=c\in[-1,1]\quad\text{a.s.}$
for such $(\mathbb{x}_{1},\mathbb{x}_{2})$ and $y$ since $\varphi$ is strictly convex. It follows that
$c=\mathbb{E}(g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2}))=f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2})\quad\text{a.s.}$
and hence
$f_{y}(\mathbb{x}_{1})-g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})=f_{y}(\mathbb{x}_{2})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})\quad\text{a.s.}$ (58)
for all $(\mathbb{x}_{1},\mathbb{x}_{2})\in G_{y}$ and $y\in N^{c}\,.$ Integrating out in (58) then gives
$f_{y}(\mathbb{x}_{1})-g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})=\int f_{y}(\mathbb{x})-g_{y}(\mathbb{x},\mathbb{z})\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}}(\mathbb{z})=0\,$,
so that $F_{Y|\mathbb{X}=\mathbb{x},\mathbb{Z}=\mathbb{z}}(y)=g_{y}(\mathbb{x},\mathbb{z})=f_{y}(\mathbb{x})=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ for $P^{(\mathbb{X},\mathbb{Z})}$-almost all $(\mathbb{x},\mathbb{z})$ and $P^{Y}$-almost every $y$. In other words, $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. Finally, a similar line of arguments yields the equivalence of $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ and conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$. $\blacksquare$
(i): Due to Theorem 1.6, we know that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies $(F_{Y}\circ q_{Y\mid\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}\mid\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$ and thus $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y\mid\mathbb{X}}^{\uparrow U}(V),U)\leq\mu(F_{Y^{\prime}}\circ q_{Y^{\prime}\mid\mathbb{X}^{\prime}}^{\uparrow U}(V),U)=\mu(Y^{\prime},\mathbb{X}^{\prime})$.
(ii): $\mathbb{X}$ and $Y$ are independent if and only if $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$ are independent; see Remark 2.8(b). Since $\mu$ characterizes independence on $\mathcal{R}^{\uparrow}$, the latter is equivalent to $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=0$.
(iii): $Y$ perfectly depends on $\mathbb{X}$ if and only if $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$ are comonotone; see Remark 2.8(b). Since $\mu$ characterizes comonotonicity on $\mathcal{R}^{\uparrow}$, the latter is equivalent to $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=1$.
(iv): By Theorem 1.4, we have $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$. Hence, (i) implies the statement.
(v): Assume that $\mu(Y,\mathbb{X})=\mu(Y,(\mathbb{X},\mathbb{Z}))$. By information monotonicity of $\preccurlyeq_{ccx}$ and due to Theorem 1.6, we know that $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. This implies $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq\mu(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)=\mu(Y,(\mathbb{X},\mathbb{Z}))=\mu(Y,\mathbb{X})$ and thus $\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=\mu(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. Since $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$, it follows that $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{c}(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$ and thus, by Theorem 1.6, $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$. However, the latter means that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. For the reverse direction, assume that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. Then $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$ and thus, by Theorem 1.6, $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. This implies $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=\mu(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)=\mu(Y,(\mathbb{X},\mathbb{Z}))$. $\blacksquare$
B.7 Proofs of Section 5
Assume that $(Y,X)\preccurlyeq_{ccx}(Y^{\prime},X^{\prime})$. Since $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$, we obtain from a variant of Theorem 1.6 that this is equivalent to
$(q_{Y|X}^{\uparrow U}(V),U)\leq_{c}(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V),U)$ (59)
for $U,V$ independent and uniform on $(0,1)$. Since $Y\uparrow_{st}X$ it follows that $P(Y\leq y\mid X=q_{X}(u))=F_{u}(y)$ and thus
$q_{Y|X}^{\uparrow U}(V)=F_{U}^{-1}(V)=F_{Y|X=q_{X}(U)}^{-1}(V),$ (60)
with $q_{Y|X}^{\uparrow u}(v)$ defined by (10). Similarly, we obtain $q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V)=F_{Y^{\prime}|X^{\prime}=q_{X^{\prime}}(U)}^{-1}(V)$. Now, observe that
$(Y,X)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y|X=q_{X}(U)}^{-1}(V),q_{X}(U))\quad\text{and}\quad(Y^{\prime},X^{\prime})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}|X^{\prime}=q_{X^{\prime}}(U)}^{-1}(V),q_{X^{\prime}}(U));$ (61)
see (38). Since the concordance order is invariant under increasing transformations, we obtain from $q_{X}=q_{X^{\prime}}$ with (59) and (60) that $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$.
To proof the reverse direction, assume that $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$. Using (61) and (60), it follows that
$(q_{Y|X}^{\uparrow U}(V),q_{X}(U))\leq_{c}(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V),q_{X^{\prime}}(U)).$ (62)
Since $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}\,$, (62) is equivalent to
$P(q_{Y|X}^{\uparrow U}(V)\leq y,U\leq F_{X}(x))\leq P(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V)\leq y,U\leq F_{X}(x))$ (63)
for all $x,y\in\mathbb{R}$; see Proposition A.1(i) and B.2. Since $q_{Y|X}^{\uparrow u}(v)$ and $q_{Y^{\prime}|X^{\prime}}^{\uparrow u}(v)$ depend on $u$ only through $q_{X}(u)$, it is not difficult to see that (63) is equivalent to $P(q_{Y|X}^{\uparrow U}(V)\leq y,U\leq u)\leq P(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V)\leq y,U\leq u)$ for all $y\in\mathbb{R}$ and $u\in(0,1)$. Using $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$, this implies $(F_{Y}\circ q_{Y|X}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V),U)$ which is by Theorem 1.6 equivalent to $(Y,X)\leq_{ccx}(Y^{\prime},X^{\prime})$. $\blacksquare$
B.8 Proofs of Appendix A
For statement (i), see e.g. [23]; for statements (ii) – (iv), see [4, Lemma A.1]. $\blacksquare$
The equivalence of (i) and (iii) is given in [4, Proposition 2.14]. The equivalence between (ii) and (iii) is a direct consequence of Proposition A.1(ii) and (iii). $\blacksquare$
The equivalence of (i) and (ii) as well as the equivalence of (iii) and (iv) follow from the invariance properties of the convex order in Lemma A.3(iv).
To show that (ii) implies (iii), we consider two cases for $v,v_{n}\in(0,1)$ with $q_{Y}(v)<q_{Y}(v_{n})$ for all $n$. In either case, we obtain for $S_{n}:=P(Y<q_{Y}(v_{n})\mid\mathbb{X})$ and $S:=P(Y\leq q_{Y}(v)\mid\mathbb{X})$ that $\mathbb{E}|S_{n}-S|=\mathbb{E}|P(q_{Y}(v)<Y<q_{Y}(v_{n})\mid\mathbb{X})|=P(q_{Y}(v)<Y<q_{Y}(v_{n}))$. In both cases, we specify $v_{n}$ such that
$\lim_{n\to\infty}\mathbb{E}|S_{n}-S|$ $=P\left(\bigcap_{n\in\mathbb{N}}\{Y<q_{Y}(v_{n})\}\right)-P(Y\leq q_{Y}(v))=0\,.$ (64)
In the first case, let $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})=\mathrm{int}(\overline{\mathsf{Ran}(F_{Y^{\prime}})})$ and $v_{n}\downarrow v$ (i.e., $v_{n}>v$ for all $n$ and $v_{n}\to v$ as $n\to\infty$). Then, $y<q_{Y}(v_{n})$ for all $n$ implies $y\leq q_{Y}(v)$ or $y\in(q_{Y}(v),q_{Y}^{+}(v))$, where $q_{Y}^{+}(t):=\sup\{x\mid F(x)\leq t\}.$ Since $\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ is a $P$-null set for this choice of $v$, we obtain (64).
In the second case, let $v\in(0,1)\setminus\overline{\mathsf{Ran}(F_{Y})}=(0,1)\setminus\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ and $v_{n}\downarrow v_{0}:=F_{Y}(q_{Y}(v))$. Then, we have that $q_{Y}(v)=q_{Y}(v_{0})$. If $F_{Y}$ is strictly increasing on $q_{Y}(v_{0}),q_{Y}(v_{0})+\varepsilon)$ for some $\varepsilon>0$, then $q_{Y}(v_{n})\downarrow q_{Y}(v_{0})$. This implies $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}(v)\geq Y\}$, which yields (64). If $F_{Y}$ is constant on $q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]$ and continuous at $q_{Y}^{+}(v_{0})$, then $q_{Y}(v_{n})\downarrow q_{Y}^{+}(v_{0})$. This implies $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}^{+}(v_{0})\geq Y\}=\{q_{Y}(v_{0})\geq Y\}\cup\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]\}$. Since $\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]\}$ is a $P$-null set, we again obtain the convergence in (64). If $F_{Y}$ is constant on $q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))$ and has a jump at $q_{Y}^{+}(v_{0})$, then $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}^{+}(v_{0})>Y\}=\{q_{Y}(v_{0})\geq Y\}\cup\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))\}$. Once again, using that $\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))\}$ is a $P$-null set, (64) follows.
The convergence in (64) implies $S_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }S$ and $\mathbb{E}|S_{n}|\to\mathbb{E}|S|$. Similar considerations yield $S^{\prime}_{n}:=P(Y^{\prime}<q_{Y^{\prime}}(v_{n})\mid\mathbb{X}^{\prime})\xrightarrow{\penalty 10000\ d\penalty 10000\ }S:=P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ and $\mathbb{E}|S^{\prime}_{n}|\to\mathbb{E}|S^{\prime}|$. Using that $S_{n}\leq_{cx}S_{n}^{\prime}$, we obtain from Lemma A.3(v) that $S\leq_{cx}S^{\prime}$. This proves (iii).
It remains to show that (iii) implies (ii). Therefore, assume that $P(Y\leq q_{Y}(u)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(u)\mid\mathbb{X}^{\prime})$ for all $u\in(0,1)\setminus N$ for some $\lambda$-null set $N\subset(0,1)$. We distinguish between two cases for $v,v_{n}\in(0,1)$ with $q_{Y}(v_{n})<q_{Y}(v)$ for all $n$. Setting $S_{n}:=P(Y\leq q_{Y}(v_{n})\mid\mathbb{X})$ and $S:=P(Y<q_{Y}(v)\mid\mathbb{X})$, we show that
$\lim_{n\to\infty}\mathbb{E}|S-S_{n}|$ $=P(Y<q_{Y}(v))-P\left(\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}\right)=0\,.$ (65)
In the first case, assume that $F_{Y}$ has no jump at $q_{Y}(v)$. Choose $v_{n}\in(0,1)\setminus N$ with $v_{n}\uparrow v$. Then, using left-continuity of $q_{Y}$, we obtain $\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}=\{Y<q_{Y}(v)\}$, which implies (65). In the second case, assume that $F_{Y}$ has a jump at $q_{Y}(v)$ and set $v_{0}:=F_{Y}^{-}(q_{Y}(v))$. The case $v_{0}=0$ is trivial. Hence, suppose $v_{0}>0$. Choose $v_{n}\in(0,1)\setminus N$ with $v_{n}\uparrow v_{0}$. Then, a distinction of cases as in the reverse direction yields $P\left(\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}\right)=P\left(\{Y<q_{Y}(v_{0})\}\right)$, which again yields (65). Both cases imply $S_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }S$ and $\mathbb{E}|S_{n}|\to\mathbb{E}|S|$. Similarly, we obtain $S_{n}^{\prime}:=P(Y^{\prime}\leq q_{Y^{\prime}}(v_{n})\mid\mathbb{X}^{\prime})\xrightarrow{\penalty 10000\ d\penalty 10000\ }P(Y^{\prime}<q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})=:S^{\prime}$ and $\mathbb{E}|S_{n}^{\prime}|\to\mathbb{E}|S^{\prime}|$. Using $S_{n}\leq_{cx}S_{n}^{\prime}$ for all $n$, Lemma A.3(v) yields $S\leq_{cx}S^{\prime}$. This proves (ii). $\blacksquare$